### v1

# Bridging the Gap Between LLMs and Human Intentions: Progresses and Challenges in Instruction Understanding, Intention Reasoning, and Reliable Generation

Abstract
Large language models (LLMs) have demonstrated exceptional capabilities in understanding and generation. However, when interacting with human instructions in real-world scenarios, LLMs still face significant challenges, particularly in accurately capturing and comprehending human instructions and intentions. This paper focuses on three challenges in LLM-based text generation tasks: instruction understanding, intention reasoning, and reliable generation. Regarding human complex instruction, LLMs have deficiencies in understanding long contexts and instructions in multi-round conversations. For intention reasoning, LLMs may have inconsistent command reasoning, difficulty reasoning about commands containing incorrect information, difficulty understanding user ambiguous language commands, and a weak understanding of user intention in commands. Besides, In terms of reliable generation, LLMs may have unstable generated content and unethical generation. To this end, we classify and analyze the performance of LLMs in challenging scenarios and conduct a comprehensive evaluation of existing solutions. Furthermore, we introduce benchmarks and categorize them based on the aforementioned three core challenges. Finally, we explore potential directions for future research to enhance the reliability and adaptability of LLMs in real-world applications.
1 Introduction
Rapid advancements with the development of large language models (LLMs) have been experienced in the field of artificial intelligence. These models, built upon massive amounts of data and extensive computing resources, have shown impressive capabilities in understanding and generating human language. Recent advancements in LLMs, including the use of scaling laws Kaplan et al. (2020), supervised fine-tuning (SFT) Wu et al. (2021), and reinforcement learning with human feedback (RLHF) Ouyang et al. (2022), have propelled these models to new heights. Researchers have explored innovative strategies like chain-of-thought reasoning (COT) Wei et al. (2022), aiming to enhance their performance in processing and generating accurate responses. However, they still struggle with more complex interactions, especially when the input data is ambiguous, incomplete, or inconsistent. Despite improvements, issues such as content hallucination Li et al. (2023) and logical misinterpretations remain prevalent. Consequently, while LLMs show promise, they are far from flawless and require further refinement to address the challenges posed by more unpredictable and complex human instructions as follows.
Figure 1: Example of large language model generation.
I. Challenge of Instruction Understanding. One of the most pressing challenges that LLMs face is instruction understanding as Figure 1(a), particularly when the user input involves complex or multi-step instructions. While models have improved in parsing relatively simple queries, they continue to encounter significant difficulties when dealing with long, context-rich instructions or when instructions are spread across multiple conversational turns. LLMs often fail to grasp subtle nuances or interpret implicit meanings embedded within the text, which leads to inaccurate or incomplete responses. Existing approaches to instruction understanding have introduced techniques like optimizing the model’s parsing abilities Teng et al. (2024), and context-aware optimization Sun et al. (2024). While these methods show promise, they often fall short when addressing the complexities and ambiguities present in real-world instructions.
II. Challenge of Intention Reasoning. Another critical area is intention reasoning as illustrated in Figure 1(b), where they struggle to align the generated responses with the user’s underlying intention. Ambiguities in language, conflicting instructions, and implicit requirements often result in models producing outputs that diverge from the user’s expectations. LLMs also face difficulties when instructions are inconsistent or contain incorrect information, which challenges the model’s ability to make accurate inferences. Various strategies, including retrieval-enhanced generation and fine-tuning techniques, have been proposed to enhance reasoning capabilities, enabling the models to better handle inconsistent or incomplete instructions. However, these methods often introduce new challenges related to bias and the inability to fully resolve conflicts in user input, further complicating the alignment between generated content and user expectations.
III. Challenge of Reliable Generation. The final major challenge is the reliable generation, which pertains to the accuracy, ethical considerations, and stability of the content they produce, such as Figure 1(c). While LLMs are generally capable of generating coherent and contextually relevant outputs, they sometimes exhibit instability, generating content that is factually incorrect, logically inconsistent, or ethically questionable. This challenge is exacerbated by the model’s inability to recognize uncertainty, which can lead to overconfident but inaccurate outputs. Recent efforts to address this issue involve techniques like uncertainty-aware fine-tuning and using external tools to evaluate output credibility. However, these approaches struggle to provide a comprehensive and reliable solution, especially in complex or dynamic contexts.
{forest}
forked edges, for tree= grow=east, reversed=true, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, anchor=center , [ Gap Between LLMs and Human Intention, ver, color=hidden-draw, fill=mygray!30, text=black [ Instruction
Understanding (§2), color=myblue!100, fill=myblue!80, text width=6em, text=black [ Long-text
Comprehension (§2.1), color=bluelight1!60, fill=bluelight1!60, text width=7em, text=black [ Information
Focusing, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ 1) Attention Sparsification Beltagy et al. (2020)
2) Attention OptimizationChen et al. (2024)
3) Position independent Training He et al. (2024a) , understanding_leaf, text width=26.5em ] ] [ Multipath
Optimization, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ 1) SFT and RL Zhang et al. (2024b)
2) Retrieval-Augmented Li et al. (2024b)
3) Recurrent Sequence Optimization Gu and Dao (2023) , understanding_leaf, text width=26.5em ] ] ] [ Multi-turn Conversation
Handling (§2.2), color=bluelight1!60, fill=bluelight1!60, text width=7em, text=black [ 1) SFT Teng et al. (2024); Sun et al. (2024), 2) RL Zhou et al. (2024) , understanding_leaf, text width=32em ] ] ] [ Intention
Reasoning (§3), color=mygreen!80, fill=mygreen!80, text width=6em, text=black [ Inconsistent Instruction
Resolution (§3.1), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Knowledge
Updating, color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ SituatedQA Zhang and Choi (2021), CDConv Zheng et al. (2022) , reasoing_leaf, text width=26.5em ] ] [ Confidence
Calibration, color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ CD2 Jin et al. (2024), MacNoise Hong et al. (2024) , reasoing_leaf, text width=26.5em ] ] ] [ Misinformation
Reasoning (§3.2), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ CKL Jang et al. (2022), StreamingQA Liska et al. (2022), BIPIA Yi et al. (2023),
CAR Weller et al. (2024) , reasoing_leaf, text width=32em ] ] [ Fuzzy Language
Interpretation (§3.3), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Folkscope Yu et al. (2023), Miko Lu et al. (2024), ATC Deng et al. (2023) , reasoing_leaf, text width=32em ] ] [ Intention
Clarification
Failure (§3.4) , color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ DeepSeek-R1 Guo et al. (2025), S1 Muennighoff et al. (2025),
SoulChat Chen et al. (2023), MoChat Mo et al. (2024) , reasoing_leaf, text width=32em ] ] ] [ Reliable
Generation (§4), color=lightkhaki!80, fill=lightkhaki!80, text width=6em, text=black [ Response Stability (§4.1), color=lightkhaki!60, fill=lightkhaki!60, text width=7em, text=black [ Fine-tuning
LLMs, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ EDL Sensoy et al. (2018), DER Amini et al. (2020),
ConformalFactuality Mohri and Hashimoto (2024), LUQ Zhang et al. (2024a) , reliable_leaf, text width=26.5em ] ] [ External Tools, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ SUE Liu et al. (2024), CalibrateMath Lin et al. (2022) , reliable_leaf, text width=26.5em ] ] ] [ Alignment (§4.2), color=lightkhaki!60, fill=lightkhaki!60, text width=7em, text=black [ Data Cleaning
and Curation, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ Stochastic Parrots Bender et al. (2021) , DeepSoftDebias Rakshit et al. , reliable_leaf, text width=26.5em ] ] [ RL-based
Alignment , color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ PPO Ouyang et al. (2022), DPO Zeng et al. (2024), GRPO Shao et al. (2024),
RLAIF Lee et al. (2024) , reliable_leaf, text width=26.5em ] ] [ In-context
Alignment, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ URIAL Lin et al. (2024), ICA Huang et al. (2024a) , reliable_leaf, text width=26.5em ] ] ] ] ]
Figure 2: Challenges and existing solutions between LLMs and Human Intentions.
Present Survey.
Facing these challenges, there is an increasing need for focused research on LLMs and their interaction with human instructions and intentions. This paper systematically analyzes LLMs’ performance in processing human instructions, highlighting three key areas: user instruction understanding, intention comprehension and reasoning, and reliable content generation. While existing review papers address model training, fine-tuning, and specific aspects of LLMs’ capabilities Lou et al. (2024); Plaat et al. (2024); Huang et al. (2024b), our focus is on the LLMs’ ability to understand and reason about user intentions. Specifically, we explore how well LLMs understand user input, reason about the user’s intention, and generate content that aligns with human intentions, minimizing hallucinations and ensuring credibility and ethical consistency.
Comparison with Previous Surveys.
While the gap between human intention and LLMs is a core challenge in generative AI, many studies focus on specific aspects of the issue, lacking a comprehensive overview. These works offer valuable insights but do not provide a systematic summary of the field. Lou et al. Lou et al. (2024) primarily address instruction following challenges in LLMs without delving into the reasoning capabilities for complex user instructions. Gao et al. analyze the four stages of human-machine LLM interaction (planning, facilitation, iteration, and testing) but overlook LLM’s understanding of user instructions. Xu et al. Xu et al. (2024b) examine the impact of various memory conflicts on LLM-generated content credibility and performance, yet do not consider reasoning or intention comprehension. Plaat et al. Plaat et al. (2024) focus on LLM Reasoning for basic mathematical problems, without exploring its applicability to broader fields. Shorinwa et al. Shorinwa et al. (2024) provide an initial analysis of LLMs uncertainty quantification, but exclude user input instructions. In contrast, our survey offers a more comprehensive perspective, with a unique classification and systematic analysis of instruction processing, addressing current solutions to key challenges.
Survey Organization.
As in Figure 2, we begin by exploring the capability of user instruction understanding (§2). Next, we analyze intention comprehension and reasoning, focusing on how models infer implicit intentions, incorporate contextual information for logical reasoning, and address inconsistencies or incomplete instructions (§3). We then examine reliable content generation, assessing the quality and credibility of model-generated outputs (§4). Following this, we review benchmarks that address these core challenges (§5). Finally, we propose potential research directions (§6) and summarize the key findings (§7).
Figure 3: Common understanding issues in LLMs include: (A) Remote Information Failure (§2.1), where the model forgets relevant information over long distances in long context (A.1). (B) Incorrect Relevance Judgment(§2.2), such as the model incorrectly associates wrong content (B.1-2.nd) from the previous turn (B.1-1.st).
2 Instruction Understanding
LLMs excel at single-turn dialogues, but struggle to understand multi-turn dialogues and long-contexts, which are commonly used by users. LLMs may forget prior information, be influenced by irrelevant data, and overlook key inputs.
2.1 Long-Text Comprehension
Challenges remain in understanding long context and this paper classifies the factors into three categories as follows:1) Information Sparsity and Redundancy. Long texts often contain redundant or irrelevant information that can obscure the task-relevant content, leading to difficulties in information extraction and potential hallucinations. 2) Remote Information Failure (Figure 3 A.1). Long contexts may cause models to forget relevant information that is distant within the text. Additionally, links between remote information across paragraphs or sentences can be difficult for models to identify, diminishing their understanding of contextual connections. 3) Attention Dilution. As context length increases, the model’s attention mechanism faces greater computational demands and struggles to assign appropriate weights to each token, making it harder to prioritize key information, particularly with complex, multi-level relationships in longer texts. This paper classifies the existing solutions into the following two categories:
Information Focusing.
Improving LLM’s ability to focus on important information in long texts involves several methods: 1) Sparsifying attention to concentrate on critical information Beltagy et al. (2020). 2) Optimizing attention to minimize redundancy and emphasize core content Chen et al. (2024). 3) Training with location-independent tasks to enhance the ability to search and react to relevant information in long contexts He et al. (2024a).
Multipath Optimization.
Various methods can enhance LLMs on long-context tasks: 1) Pre-training with extended context windows and reinforcement learning for fine-tuning to optimize long-context understanding Zhang et al. (2024b). 2) Combining retrieval-based models with generative models on long-context tasks Li et al. (2024b). 3) Leveraging cyclic sequence models’ linear scaling property for better inference efficiency Gu and Dao (2023).
2.2 Multi-Turn Conversation Handling
This paper categorizes the challenges faced by existing LLMs when understanding multi-turn conversations into three categories, as follows: 1) Capability Weakening. Current supervised instruction fine-tuning (SIFT) and RLHF may even impair multi-turn capabilitiesWang et al. (2024), with models struggling on complex reasoning tasks that span multiple rounds, such as those requiring evidence collection and conclusions Banatt et al. (2024). Additionally, multi-turn dialogs increase the vulnerability of LLMs to adversarial attacks, where malicious users can mask harmful intentions across multiple rounds, leading to the generation of misleading or harmful content Agarwal et al. (2024). 2) Error Propagation. Instruction comprehension errors accumulate across rounds, leading to an escalating failure rate in subsequent responses He et al. (2024b), which may snowball into larger issues such as biased or incorrect outputs Fan et al. (2024). 3) Incorrect Relevance Judgment (Figure 3 B.1) LLMs often struggle to identify relevant content in multi-turn dialogs, failing to properly link content from previous rounds or to discern ellipsis and implicit meaning inherent in user commands Sun et al. (2024).
To solve above challenges, This paper categorizes existing solutions into two types: supervised fine-tuning methods using multi-turn dialogue data, enhanced by techniques like optimized instruction parsing Teng et al. (2024) and context-aware preference strategies Sun et al. (2024); and reinforcement learning methods tailored for multi-turn dialogue, with improvements such as hierarchical reinforcement learning Zhou et al. (2024).
3 Intention Reasoning
User instructions often lack clarity due to language ambiguities. While humans can infer intention, LLMs struggle with misinterpreting ambiguous inputs, leading to errors. We explore causes and solutions for intention errors, focusing on inconsistent instructions, misinformation, fuzzy language, and intention clarification.
Figure 4: Common reasoning issues in LLMs include: (A) Inconsistent Instruction Reasoning (§3.1), where LLMs fail to detect conflicting inputs (A.1) or overlooking logical inconsistencies (A.2), (B) Misinformation Reasoning (§3.2), caused by temporal misalignment leading to outdated responses (B.1) or data contamination resulting in misleading outputs (B.2). (C) Fuzzy Language Interpretation (§3.3), where the model relies on biases for fuzzy queries (C.1) or defaults to a response without seeking clarification (C.2). and (D) Intention Clarification Failure (§3.4) , where it misinterprets sarcasm (D.1) or ignores prior emotional context (D.2).
3.1 Inconsistent Instruction Reasoning
In natural language communication, humans easily identify inconsistencies using context and prior knowledge, whereas LLMs struggle, often accept contradictory inputs, and generate unreliable answers. This phenomenon has been observed across multiple question-answering generation tasks Li et al. (2024a); Zheng et al. (2022), and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Ignoring input errors (Figure 4 A.1). The model ignores the input errors and gives an answer, resulting in the model assigning the same weight to each context given by the user, which in turn affects the generation of the answer. 2) Inability to detect user inconsistencies (Figure 4 A.2). In the premise that the model has learned the knowledge, the model still has difficulty detecting user inconsistencies. To address inconsistent instruction reasoning issues, existing solutions primarily adopt the following two approaches:
Knowledge Updating.
SituatedQA Zhang and Choi (2021) attempts to enhance model performance by updating the knowledge base. Additionally, CDConv Zheng et al. (2022) simulates common user behaviors to trigger chatbots through an automated dialogue generation method, generating contradictions for training purposes.
Confidence Calibration.
Given the high cost of data annotation and model fine-tuning, some researchers have sought alternative approaches by introducing additional processing techniques. CD2 Jin et al. (2024) maximizes probabilistic output and calibrates model confidence under knowledge conflicts using conflict decoupling and comparison decoding methods. MacNoise Hong et al. (2024) enhances contradiction retrieval in an augmented generative system by explicitly fine-tuning a discriminator or prompting LLMs to improve contradiction discrimination.
3.2 Misinformation Reasoning
Erroneous instructions mislead model outputs more severely than inconsistent ones, as they lack obvious contradictions, requiring the model to comprehend, reason, compare input knowledge with its parameterized knowledge, and make objective judgments Cheang et al. (2023); Xu et al. (2024a). From the input perspective, this paper classifies the sources of erroneous information into two categories as follows: 1) Temporal Alignment Failure (Figure 4 B.1). arises when the knowledge provided by the user and the model is temporally misaligned due to updates occurring at different times, leading to inconsistent responses. Such discrepancies typically originate during the training process. 2) Information Contamination (Figure 4 B.2). refers to the degradation of model quality caused by the intentional distortion of input data.
To solve the above problems, existing methods mainly focus on improving model susceptibility in the face of internal and external knowledge conflicts through targeted fine-tuning and processing. CKL Jang et al. (2022) ensures that the model’s knowledge is updated in a timely manner through an online approach, although this approach is slightly weaker than re-training in terms of effectiveness Liska et al. (2022). BIPIA Yi et al. (2023) used adversarial training to combat the effects of information pollution and improve model robustness. CAR Weller et al. (2024) achieved nearly 20% improvement by discriminating external knowledge that may not be contaminated in the RAG system.
3.3 Fuzzy Language Interpretation
When user instructions contain fuzzy terms (e.g., polysemy or vagueness), LLMs may select an incorrect interpretation from multiple possibilities, potentially leading to misleading responses. This phenomenon has been observed across multiple information-seeking tasks Kim et al. (2024), and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Self-defined problem (Figure 4 C.1). When the user inputs content with fuzzy sentences, LLMs may choose to generate content based on the preferences of its own training data. 2) Select data based on fuzzy input (Figure 4 C.2). In response to ambiguous user input, LLMs may select a default explanation without actively asking the user to clarify.
To solve the problem, researchers have started to experiment with cue engineering and relation alignment. Folkscope Yu et al. (2023) proposed the FolkScope framework, which uses a large language model to analyze and discriminate users’ fuzzy purchasing intention. Miko Lu et al. (2024) introduces a multimodal hierarchical intention generation framework that interprets users’ posting behaviors by analyzing the fuzzy information they share on social platforms. ATC Deng et al. (2023) utilizes the Active Thinking Chain cueing scheme, which enhances the proactivity of a biglanguage model by adding goal-planning capabilities to the descriptive reasoning chain.
3.4 Intention Clarification Failure
Unlike humans, who reason based on experience, LLMs lack real-world common sense and thus struggle to infer complex contexts beyond their input data. Moreover, they often fail to maintain a consistent reasoning trajectory across long texts, complex contexts, or multi-turn conversations. When handling intricate intentions or sentiment shifts, LLMs may struggle to retain prior context, leading to errors in inferring implicit user needs. We categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Fails to detect sarcasm (Figure 4 D.1), when LLMs fails to understand the sarcastic intention of the user. 2) Ignores prior emotional context (Figure 4 D.2), when LLMs focus only on the second half of the sentence and ignore the emotions in the previous round of dialog.
To solve above problems, researchers have started to try to construct multi-domain datasets containing implicit intentions to strengthen the ability of the LLMs to reason about complex intentions and user emotions in multi-round interaction scenarios. DeepSeek-R1 Guo et al. (2025) enhances its understanding of human intention through a structured process with two RL stages for refining reasoning patterns and aligning with human preferences. S1 Muennighoff et al. (2025) uses budget forcing to control the number of thinking tokens. The upper limit is terminated early by a delimiter, while the lower limit prohibits delimiters and adds ”wait” to guide in-depth reasoning and optimize the quality of the answer. SoulChat Chen et al. (2023) fine-tuned LLMs by constructing a dataset containing more than 2 million samples of multi-round empathic conversations. MoChat Mo et al. (2024) constructs multi-round dialogues for spatial localization by using joint grouping spatio-temporal localization.
4 Realiable Generation
Despite strong performance, LLMs struggle with output reliability. Trained on large corpora using maximum likelihood estimation, they generate deterministic responses. While effective on familiar data, they often produce unstable or incomplete responses to unseen inputs, undermining reliability. We analyze LLM response stability and alignment in depth.
4.1 Response Stability
The knowledge acquired by the LLMs is generally determined in the pre-training stage and stored in a parameterized form. For data in a specific field, the current model is generally optimized by fine-tuning instructions so that it outputs what humans want Radford et al. (2019). If knowledge samples that the model has not seen are used in instruction-tuning, it will inevitably cause the model to give a definite response to unknown inputs, and there is a high probability that an answer will be fabricated. This is the over-confidence of the model that causes the model to output unreliable answers, and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Fabricated incorrect information (Figure 5 A.1). When the model’s knowledge did not match the input question, it fabricated information that did not match the facts. 2) Incorrect context output (Figure 5 A.2). When the model’s knowledge did match the input question, it output incorrect context information. To address these issues, researchers have explored uncertainty, which quantifies the credibility and stability of model outputs as follows.
Figure 5: Common generation issues in LLMs include: (A) Unstable Content Generation (§4.1), where the model fabricates details when it lacks relevant knowledge (A.1) or produces incorrect contextual information despite having relevant knowledge (A.2). (B) Misalignment with Human Values (§4.2), where the model generates harmful or offensive content (B.1) or provides responses that conflict with moral and ethical standards (B.2).
Fine-tuning LLMs.
To make LLMs more accurate in estimating uncertainty, existing methods fine-tune models Sensoy et al. (2018); Amini et al. (2020). LUQ Zhang et al. (2024a) is a novel sampling-based uncertainty quantification method specifically designed for long texts. ConformalFactuality Mohri and Hashimoto (2024) defines the associated uncertainties for each possible output.
External Tools.
Fine-tuning LLMs typically demands substantial computing resources and slow training; therefore, reducing computational overhead is crucial for improving efficiency. Researcher has proposed methods to evaluate the uncertainty of model outputs through external tools Liu et al. (2024). CalibrateMath Lin et al. (2022) assesses uncertainty by requiring models to generate numerical answers with confidence levels, evaluating their reliability.
4.2 Alignment
Despite the impressive capabilities of large language models (LLMs), they have raised significant concerns regarding the unsafe or harmful content they may generate. LLMs are typically trained on vast datasets scraped from the internet, including inappropriate or harmful content Bender et al. (2021). This means that the models may inadvertently produce outputs misaligned with human values as follows: 1) Generation of Toxic Content (Figure 5 B.1). LLMs may generate toxic content, such as hate speech or offensive comments, when asked to respond to sensitive topics Luong et al. (2024); Dutta et al. (2024). 2) Conflicts with Moral/Ethical Standards (Figure 5 B.2). LLMs might produce outputs that conflict with moral or ethical standards, such as guiding illegal activities Ramezani and Xu (2023); Abdulhai et al. (2024). To tackle the previously mentioned concerns regarding unsafe or harmful content produced by LLMs, researchers have focused on various stages:
Pretraining Data Cleaning and Curation.
To minimize the risks associated with harmful or inappropriate content, LLM training datasets should undergo rigorous cleaning processes Bender et al. (2021), such as filtering out toxic language, hate speech, and harmful stereotypes. Tools like word embedding debiasing methods Rakshit et al. can help identify and remove toxic and biased content.
Table 1: A selection of widely used benchmark datasets for evaluating LLMs. The ’Type’ column categorizes the benchmarks based on task types, while ’Lang.’ indicates the language of the dataset, and ’Citation’ denotes the number of citations each benchmark has received.
Type Benchmark Year Lang. Description Citation
[…]
MT-BENCH-101 Bai et al. (2024) 2024 En Evaluation of multi-turn dialogue capability in LLMs. 39
LLM Reasoning BIPIA Yi et al. (2023) 2023 En Assesses LLMs vulnerability to hint injection attacks. 60
Miko Lu et al. (2024) 2024 En Evaluates multimodal LLMs in understanding social intention. 2
CONTRADOC Li et al. (2024a) 2023 En Detects self-contradictions in long documents. 7
CDCONV Zheng et al. (2022) 2022 Zh Detects contradictions in Chinese dialogues. 12
LLM Generation Open-LLM-Leaderboard Ye et al. (2024) 2024 En Assesses LLMs uncertainty in generated content. 19
ETHICS Dataset Hendrycks et al. (2021) 2020 En Evaluates LLMs on moral reasoning and ethics. 468
FACTOR Muhlgay et al. (2024) 2023 En Measures factual accuracy in LLM-generated text. 71
Reinforcement Learning-based Alignment.
To further align LLMs with human values and societal norms, reinforcement learning methods, such as RLHF and its advanced variants, such as PPO Ouyang et al. (2022), DPO Zeng et al. (2024), and GRPO Shao et al. (2024), are very essential. As an extension of RLHF, RLAIF Lee et al. (2024) leverages AI systems to assist in the feedback process, making evaluation and fine-tuning more scalable.
In-context Alignment.
In-context alignment leverages the ability of LLMs to adapt their responses based on a few examples provided in the prompt. Lin et al. (2024) demonstrates that effective alignment can be achieved purely through ICL with just a few stylistic examples and a system prompt. Huang et al. (2024a) explored the effectiveness of different components of In-context alignment, and found that examples within the context are crucial for enhancing alignment capabilities.
5 Benchmark
This section covers benchmarks for LLMs in instruction understanding, reasoning, and generation (Table 1).
5.1 Benchmarking Instruction Understanding
Instruction understanding in LLMs involves extracting key information, maintaining coherence, and adapting to dynamic conversation changes, especially in long or multi-round dialogues. Zhang et al. Zhang et al. (2024c) provides a benchmark consisting of 12 tasks, with an average data length exceeding 100K tokens, designed to evaluate LLMs’ ability to handle long contexts. The results show that the performance of LLMs significantly declines when dealing with long contexts, suggesting further improvements. MT-BENCH-101 Bai et al. (2024) constructs a three-layer hierarchical capability classification and 1,388 dialogue pairs across 13 tasks, providing a comprehensive evaluation of LLMs’ fine-grained capabilities in multi-turn dialogue. Experimental results show that commonly used techniques do not significantly enhance multi-turn capabilities.
5.2 Benchmarking LLM Reasoning
LLM Reasoning involves inferring user intentions by interpreting both explicit and implicit language cues. BIPIA Yi et al. (2023) evaluates LLMs under indirect hint injection attacks across five scenarios and 250 targets, revealing vulnerabilities in all models, with GPT-3.5-turbo and GPT-4 exhibiting notably higher susceptibilities. Miko Lu et al. (2024) assesses multimodal models in understanding social media user intentions. The benchmark, which includes 979 social media entries, shows that multimodal LLMs outperform text-only models like LLama2-7B and GLM4. Incorporating image data enhances the model’s ability to interpret user intentions, improving accuracy. CONTRADOC Li et al. (2024a) is the first dataset for analyzing self-contradictions in long documents. Evaluation of GPT-4, PaLM2, and other LLMs on this dataset reveals that while GPT-4 outperforms others and even surpasses human performance, it still struggles with complex contradictions requiring nuanced reasoning. CDCONV Zheng et al. (2022) focuses on contradiction detection in Chinese dialogues, containing over 12,000 dialogue rounds. The study shows that the Hierarchical method consistently outperforms others in detecting contradictions, highlighting the importance of accurate contextual modeling in dialogue understanding.
5.3 Benchmarking LLM Generation
LLM Generation assesses a model’s ability to understand user instructions, avoid fabricating false information, and generate accurate, contextually appropriate responses. Open-LLM-Leaderboard Ye et al. (2024) introduces a novel benchmark that integrates uncertainty quantification to evaluate the reliability of content generation across tasks like QA, comprehension, and dialogue. Results show that larger models often exhibit greater uncertainty, and fine-tuned models tend to have higher uncertainty despite higher accuracy. ETHICS Hendrycks et al. (2021) evaluates whether generated content aligns with human ethical values, such as justice and well-being. The study finds that while models like GPT-3 show promise in predicting human moral judgments, they still need improvement in this domain. FACTOR Muhlgay et al. (2024) addresses the evaluation of factuality in LLMs by providing a scalable method that ensures diverse and rare facts are considered. Testing models such as GPT-2 and GPT-Neo show that, while benchmark scores correlate with perplexity, they better reflect factuality in open-ended generation, especially when retrieval augmentation is applied.
6 Future Directions
This section summarizes ongoing challenges in instruction understanding, reasoning, and reliable generation with LLMs and outlines potential future research directions.
Automated Annotation Framework.
Although LLMs excel in general-domain tasks, they often produce hallucinated or incomplete content in specialized fields due to limited domain-specific training data. While contextual learning and instruction fine-tuning methods have been explored to address this issue, manual data annotation remains labor-intensive and prone to quality inconsistencies. An automated annotation framework could streamline data labeling, enhancing model performance in specialized fields by ensuring higher quality and scalability of domain-specific training datasets.
GraphRAG.
LLMs have shown impressive language generation capabilities through pre-training on large datasets, but their reliance on static data often results in inaccurate or fictional content, particularly in domain-specific tasks. The Graph-enhanced generation approach aims to tackle this by leveraging KGs and GNNs for precise knowledge retrieval. Despite its advantages, GraphRAG faces challenges in capturing structural information during graph reasoning tasks and struggles with multi-hop retrieval accuracy and conflict resolution between external and internal knowledge. Future work should focus on refining retrieval strategies and improving the stability and accuracy of GraphRAG in complex tasks.
Quantifying Uncertainty in LLMs.
As LLMs like LLama and ChatGPT have revolutionized content generation, ensuring the reliability of their outputs remains a critical challenge. Uncertainty quantification is a promising approach to address this, enabling models to provide a confidence assessment alongside their responses. Evidence learning Sensoy et al. (2018); Amini et al. (2020), an emerging method in uncertainty representation, offers a reliable approach for quantifying uncertainty directly from data. However, its application to LLMs is computationally intensive, and most existing work focuses on small-scale models. Future research should aim to optimize uncertainty quantification for large-scale models efficiently.
Balancing Safety and Performance.
Although advancements in alignment techniques have improved factual accuracy and safety, they often come at the cost of the model’s creativity and fluency. Striking a balance between safety and performance is crucial. Future research should explore new alignment methods that ensure both the safety and usability of LLMs, optimizing the trade-off between generating reliable, safe content and maintaining the model’s creative and contextual capabilities.
7 Conclusion
This paper analyzes LLMs’ performance in processing user instructions. Despite progress in natural language understanding, LLMs struggle with complex, inconsistent instructions, often resulting in biases, errors, and hallucinations. Improvements through prompt engineering, knowledge retrieval, model expansion, fine-tuning, and RLHF have not fully addressed LLMs’ limitations in reasoning and comprehension, limiting their real-world applicability. We identify three key bottlenecks: instruction comprehension, reasoning ability, and reliable content generation. Future research should focus on enhancing reasoning for complex instructions and aligning outputs with user intent, to improve LLMs’ adaptability and reliability in real-world tasks.

### v2

@@
-Bridging the Gap Between LLMs and Human Intentions: Progresses and Challenges in Instruction Understanding, Intention Reasoning, and Reliable Generation
+A Survey of Progress and Challenges in Aligning LLMs with Human Intentions
 Abstract
@@
 However, when interacting with human instructions in real-world scenarios, LLMs still face significant challenges, particularly in accurately capturing and comprehending human instructions and intentions.
-This paper focuses on three challenges in LLM-based text generation tasks: instruction understanding, intention reasoning, and reliable generation.
+This paper focuses on three challenges in LLM-based text generation tasks: instruction understanding, intention reasoning, and Reliable Dialog Generation.
 Regarding human complex instruction, LLMs have deficiencies in understanding long contexts and instructions in multi-round conversations.
 For intention reasoning, LLMs may have inconsistent command reasoning, difficulty reasoning about commands containing incorrect information, difficulty understanding user ambiguous language commands, and a weak understanding of user intention in commands.
-Besides, In terms of reliable generation, LLMs may have unstable generated content and unethical generation.
+Besides, In terms of Reliable Dialog Generation, LLMs may have unstable generated content and unethical generation.
 To this end, we classify and analyze the performance of LLMs in challenging scenarios and conduct a comprehensive evaluation of existing solutions.
@@
 Finally, we explore potential directions for future research to enhance the reliability and adaptability of LLMs in real-world applications.
+A Survey of Progress and Challenges in Aligning LLMs
+with Human Intentions
+Zongyu Chang1thanks: These authors contributed equally to this work., Feihong Lu2*, Ziqin Zhu2, Qian Li1, Cheng Ji2, Zhuo Chen1, Hao Peng2, Yang Liu5, Ruifeng Xu3, Yangqiu Song4, Jianxin Li2, Shangguang Wang1 1Beijing University of Posts and Telecommunications, 2Beihang University, 3Harbin Institute of Technology, Shenzhen, 4Hong Kong University of Science and Technology, 5Chinese Academy of Sciences {changzongyu, li.qian, sgwang}@bupt.edu.cn, {lufeihong, zhuziqin, jicheng, lijx}@buaa.edu.cn, xuruifeng@hit.edu.cn, yqsong@cse.ust.hk
 1 Introduction
@@
 Consequently, while LLMs show promise, they are far from flawless and require further refinement to address the challenges posed by more unpredictable and complex human instructions as follows.
-Figure 1: Example of large language model generation.
+Figure 1: Example of LLMs generation.
+Figure 2: Unlike previous surveys on LLMs, we do not consider the alignment between LLMs and humans as an isolated process,instead, we view it as a continuous and dynamic information processing process consisting of instruction understanding, intention reasoning, and reliable dialogue generation.
 I.
 Challenge of Instruction Understanding.
-One of the most pressing challenges that LLMs face is instruction understanding as Figure 1(a), particularly when the user input involves complex or multi-step instructions.
+One of the most pressing challenges that LLMs face is instruction understanding as Figure 1(a) and Figure (2I), particularly when the user input involves complex or multi-step instructions.
 While models have improved in parsing relatively simple queries, they continue to encounter significant difficulties when dealing with long, context-rich instructions or when instructions are spread across multiple conversational turns.
@@
 (2024).
-While these methods show promise, they often fall short when addressing the complexities and ambiguities present in real-world instructions.
+While these methods show promise, they often fall short when addressing the complexities and ambiguities present in instructions.
 II.
 Challenge of Intention Reasoning.
-Another critical area is intention reasoning as illustrated in Figure 1(b), where they struggle to align the generated responses with the user’s underlying intention.
+Another critical area is intention reasoning as illustrated in Figure 1(b) and Figure 2 (II), where they struggle to align the generated responses with the user’s underlying intention.
 Ambiguities in language, conflicting instructions, and implicit requirements often result in models producing outputs that diverge from the user’s expectations.
@@
 III.
-Challenge of Reliable Generation.
-The final major challenge is the reliable generation, which pertains to the accuracy, ethical considerations, and stability of the content they produce, such as Figure 1(c).
+Challenge of Reliable Dialog Generation.
+The final major challenge is the reliable dialog generation, which pertains to the accuracy, ethical considerations, and stability of the content they produce, such as Figure 1(c) and Figure 2 (III.
 While LLMs are generally capable of generating coherent and contextually relevant outputs, they sometimes exhibit instability, generating content that is factually incorrect, logically inconsistent, or ethically questionable.
@@
 Comprehension (§2.1), color=bluelight1!60, fill=bluelight1!60, text width=7em, text=black [ Information
-Focusing, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ 1) Attention Sparsification Beltagy et al. (2020)
-2) Attention OptimizationChen et al. (2024)
+Focusing, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ 1) Attention Sparsification Beltagy et al. (2020),
+2) Attention Optimization Chen et al. (2024),
 3) Position independent Training He et al. (2024a) , understanding_leaf, text width=26.5em ] ] [ Multipath
@@
 3) Recurrent Sequence Optimization Gu and Dao (2023) , understanding_leaf, text width=26.5em ] ] ] [ Multi-turn Conversation
-Handling (§2.2), color=bluelight1!60, fill=bluelight1!60, text width=7em, text=black [ 1) SFT Teng et al.
-(2024); Sun et al.
-(2024), 2) RL Zhou et al.
-(2024) , understanding_leaf, text width=32em ] ] ] [ Intention
+Handling (§2.2), color=bluelight1!60, fill=bluelight1!60, text width=7em, text=black [ Soft
+Fine-tuning, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ Multi-turn SFT Teng et al. (2024), Parrot Sun et al. (2024) , understanding_leaf, text width=26.5em ] ] [ Reforcement
+Learning, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ ArCHer Zhou et al. (2024) , understanding_leaf, text width=26.5em ] ] ] ] [ Intention
 Reasoning (§3), color=mygreen!80, fill=mygreen!80, text width=6em, text=black [ Inconsistent Instruction
@@
 Calibration, color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ CD2 Jin et al. (2024), MacNoise Hong et al. (2024) , reasoing_leaf, text width=26.5em ] ] ] [ Misinformation
-Reasoning (§3.2), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ CKL Jang et al. (2022), StreamingQA Liska et al. (2022), BIPIA Yi et al. (2023),
-CAR Weller et al. (2024) , reasoing_leaf, text width=32em ] ] [ Fuzzy Language
-Interpretation (§3.3), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Folkscope Yu et al.
+Reasoning (§3.2), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Targeted
+Fine-tuning, color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ CKL Jang et al. (2022), StreamingQA Liska et al. (2022), BIPIA Yi et al. (2025),
+KC-LLMs Wang et al. (2023), CAR Weller et al. (2024) , reasoing_leaf, text width=26.5em ] ] ] [ Fuzzy Language
+Interpretation (§3.3), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Clue
+Engineering , color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ Folkscope Yu et al.
 (2023), Miko Lu et al.
 (2024), ATC Deng et al.
-(2023) , reasoing_leaf, text width=32em ] ] [ Intention
+(2023) , reasoing_leaf, text width=26.5em ] ] ] [ Intention
 Clarification
-Failure (§3.4) , color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ DeepSeek-R1 Guo et al. (2025), S1 Muennighoff et al. (2025),
-SoulChat Chen et al. (2023), MoChat Mo et al. (2024) , reasoing_leaf, text width=32em ] ] ] [ Reliable
+Failure (§3.4) , color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Deep
+Reasoning , color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ DeepSeek-R1 Guo et al. (2025), S1 Muennighoff et al. (2025),
+SoulChat Chen et al. (2023), MoChat Mo et al. (2024) , reasoing_leaf, text width=26.5em ] ] ] ] [ Reliable
 Generation (§4), color=lightkhaki!80, fill=lightkhaki!80, text width=6em, text=black [ Response Stability (§4.1), color=lightkhaki!60, fill=lightkhaki!60, text width=7em, text=black [ Fine-tuning
@@
 (2024), CalibrateMath Lin et al.
-(2022) , reliable_leaf, text width=26.5em ] ] ] [ Alignment (§4.2), color=lightkhaki!60, fill=lightkhaki!60, text width=7em, text=black [ Data Cleaning
+, reliable_leaf, text width=26.5em ] ] ] [ Alignment (§4.2), color=lightkhaki!60, fill=lightkhaki!60, text width=7em, text=black [ Data Cleaning
 and Curation, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ Stochastic Parrots Bender et al.
@@
 Facing these challenges, there is an increasing need for focused research on LLMs and their interaction with human instructions and intentions.
-This paper systematically analyzes LLMs’ performance in processing human instructions, highlighting three key areas: user instruction understanding, intention comprehension and reasoning, and reliable content generation.
+This paper systematically analyzes LLMs’ performance in processing human instructions, highlighting three key areas: user instruction understanding, intention comprehension and reasoning, and reliable dialog generation.
 While existing review papers address model training, fine-tuning, and specific aspects of LLMs’ capabilities Lou et al.
@@
 (2024b), our focus is on the LLMs’ ability to understand and reason about user intentions.
-Specifically, we explore how well LLMs understand user input, reason about the user’s intention, and generate content that aligns with human intentions, minimizing hallucinations and ensuring credibility and ethical consistency.
+Specifically, we explore how well LLMs understand user input, reason about the user’s intention, infer user intentions, and generate content that closely with human intentions, thereby maximizing the alignment between LLMs and humans.
 Comparison with Previous Surveys.
@@
 Xu et al.
-(2024b) examine the impact of various memory conflicts on LLM-generated content credibility and performance, yet do not consider reasoning or intention comprehension.
+examine the impact of various memory conflicts on LLM-generated content credibility and performance, yet do not consider reasoning or intention comprehension.
 Plaat et al.
@@
 (2024) provide an initial analysis of LLMs uncertainty quantification, but exclude user input instructions.
-In contrast, our survey offers a more comprehensive perspective, with a unique classification and systematic analysis of instruction processing, addressing current solutions to key challenges.
+In contrast, our survey offers a more comprehensive perspective, as shown in Figure 2 and Figure 3, with a unique classification and systematic analysis of instruction processing, while addressing current solutions to key challenges.
 Survey Organization.
 As in Figure 3, we begin by exploring the capability of user instruction understanding (§2).
-Next, we analyze intention comprehension and reasoning, focusing on how models infer implicit intentions, incorporate contextual information for logical reasoning, and address inconsistencies or incomplete instructions (§3).
-We then examine reliable content generation, assessing the quality and credibility of model-generated outputs (§4).
-Following this, we review benchmarks that address these core challenges (§5).
+Next, we focu on how models infer implicit intentions, incorporate contextual information for logical reasoning, and address inconsistencies or incomplete instructions (§3).
+We then examine reliable dialog generation, assessing the quality and credibility of model-generated outputs (§4).
+Next, we briefly analyze the problems faced by LLMs in face of different challenges(§5) and review the benchmarks(§6) for the above problems.
 Finally, we propose potential research directions (§7) and summarize the key findings (§8).
-Figure 3: Common understanding issues in LLMs include: (A) Remote Information Failure (§2.1), where the model forgets relevant information over long distances in long context (A.1).
-(B) Incorrect Relevance Judgment(§2.2), such as the model incorrectly associates wrong content (B.1-2.nd) from the previous turn (B.1-1.st).
 2 Instruction Understanding
@@
 2.1 Long-Text Comprehension
-Challenges remain in understanding long context and this paper classifies the factors into three categories as follows:1) Information Sparsity and Redundancy.
-Long texts often contain redundant or irrelevant information that can obscure the task-relevant content, leading to difficulties in information extraction and potential hallucinations.
-2) Remote Information Failure (Figure 3 A.1).
+Understanding lengthy textual instructions remains a significant hurdle for large language models (LLMs), as real-world human instructions are often expressed in loose, unstructured natural language, contrasting with the explicitly defined tasks and structured labeling commonly employed in cue word engineering, so we categorize the relevant factors into the following three categories:1) Information Sparsity and Redundancy.
+Long texts often contain redundant or irrelevant information that can obscure the task-relevant content, leading to difficulties in information extraction.
+2) Remote Information Failure (Figure 4).
 Long contexts may cause models to forget relevant information that is distant within the text.
@@
 2.2 Multi-Turn Conversation Handling
-This paper categorizes the challenges faced by existing LLMs when understanding multi-turn conversations into three categories, as follows: 1) Capability Weakening.
+Multi-turn conversation serves as a fundamental interaction mode between LLMs and humans.
+Given the challenges users face in providing complete and precise instructions in a single turn, they often opt to refine and clarify their intentions incrementally through iterative exchanges.
+However, due to the characteristics of real-world conversations, such as the constant changes in user intentions and long-distance dependencies, LLMs still faces significant challenges in achieving coordinated multi-round interactions with humans.This paper categorizes the challenges faced by existing LLMs when understanding multi-turn conversations into three categories, as follows: 1) Capability Weakening.
 Current supervised instruction fine-tuning (SIFT) and RLHF may even impair multi-turn capabilitiesWang et al.
@@
 (2024).
-3) Incorrect Relevance Judgment (Figure 3 B.1) LLMs often struggle to identify relevant content in multi-turn dialogs, failing to properly link content from previous rounds or to discern ellipsis and implicit meaning inherent in user commands Sun et al.
+3) Incorrect Relevance Judgment (Figure 5 Q.1) LLMs often struggle to identify relevant content in multi-turn dialogs, failing to properly link content from previous rounds or to discern ellipsis and implicit meaning inherent in user commands Sun et al.
 (2024).
@@
 We explore causes and solutions for intention errors, focusing on inconsistent instructions, misinformation, fuzzy language, and intention clarification.
-Figure 4: Common reasoning issues in LLMs include: (A) Inconsistent Instruction Reasoning (§3.1), where LLMs fail to detect conflicting inputs (A.1) or overlooking logical inconsistencies (A.2), (B) Misinformation Reasoning (§3.2), caused by temporal misalignment leading to outdated responses (B.1) or data contamination resulting in misleading outputs (B.2).
-(C) Fuzzy Language Interpretation (§3.3), where the model relies on biases for fuzzy queries (C.1) or defaults to a response without seeking clarification (C.2).
-and (D) Intention Clarification Failure (§3.4) , where it misinterprets sarcasm (D.1) or ignores prior emotional context (D.2).
 3.1 Inconsistent Instruction Reasoning
@@
 (2024a); Zheng et al.
-(2022), and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Ignoring input errors (Figure 4 A.1).
+(2022), and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Ignoring input errors (Figure 6 Q.1).
 The model ignores the input errors and gives an answer, resulting in the model assigning the same weight to each context given by the user, which in turn affects the generation of the answer.
-2) Inability to detect user inconsistencies (Figure 4 A.2).
+2) Inability to detect user inconsistencies (Figure 6 Q.2).
 In the premise that the model has learned the knowledge, the model still has difficulty detecting user inconsistencies.
@@
 (2024) maximizes probabilistic output and calibrates model confidence under knowledge conflicts using conflict decoupling and comparison decoding methods.
-MacNoise Hong et al.
-(2024) enhances contradiction retrieval in an augmented generative system by explicitly fine-tuning a discriminator or prompting LLMs to improve contradiction discrimination.
 3.2 Misinformation Reasoning
@@
 (2023); Xu et al.
-(2024a).
-From the input perspective, this paper classifies the sources of erroneous information into two categories as follows: 1) Temporal Alignment Failure (Figure 4 B.1).
+(2024).
+From the input perspective, this paper classifies the sources of erroneous information into two categories as follows: 1) Temporal Alignment Failure (Figure 7 Q.1).
 arises when the knowledge provided by the user and the model is temporally misaligned due to updates occurring at different times, leading to inconsistent responses.
 Such discrepancies typically originate during the training process.
-2) Information Contamination (Figure 4 B.2).
+2) Information Contamination (Figure 7 Q.2).
 refers to the degradation of model quality caused by the intentional distortion of input data.
@@
 (2022).
+RKC-LLMs Wang et al.
+(2023) allows a large model to recognize knowledge conflicts by means of instructional fine-tuning that identifying specific passages of conflicting information.
 BIPIA Yi et al.
@@
 This phenomenon has been observed across multiple information-seeking tasks Kim et al.
-(2024), and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Self-defined problem (Figure 4 C.1).
+(2024), and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Self-defined problem (Figure 8 Q.1).
 When the user inputs content with fuzzy sentences, LLMs may choose to generate content based on the preferences of its own training data.
-2) Select data based on fuzzy input (Figure 4 C.2).
+2) Select data based on fuzzy input (Figure 8 Q.2).
 In response to ambiguous user input, LLMs may select a default explanation without actively asking the user to clarify.
-To solve the problem, researchers have started to experiment with cue engineering and relation alignment.
+To solve this problem, researchers have started to parse the semantic information expressed by users through clue engineering.
 Folkscope Yu et al.
@@
 Miko Lu et al.
-(2024) introduces a multimodal hierarchical intention generation framework that interprets users’ posting behaviors by analyzing the fuzzy information they share on social platforms.
+(2024) introduces a hierarchical intention generation framework that interprets users’ posting behaviors by analyzing the fuzzy information they share on social platforms.
 ATC Deng et al.
@@
 When handling intricate intentions or sentiment shifts, LLMs may struggle to retain prior context, leading to errors in inferring implicit user needs.
-We categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Fails to detect sarcasm (Figure 4 D.1), when LLMs fails to understand the sarcastic intention of the user.
-2) Ignores prior emotional context (Figure 4 D.2), when LLMs focus only on the second half of the sentence and ignore the emotions in the previous round of dialog.
-To solve above problems, researchers have started to try to construct multi-domain datasets containing implicit intentions to strengthen the ability of the LLMs to reason about complex intentions and user emotions in multi-round interaction scenarios.
+We categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Fails to detect sarcasm (Figure 9 Q.1), when LLMs fails to understand the sarcastic intention of the user.
+2) Ignores prior emotional context (Figure 9 Q.2), when LLMs focus only on the second half of the sentence and ignore the emotions in the previous round of dialog.
+To solve above problems, researchers have started to try to construct multi-domain datasets Chen et al.
+(2023) containing implicit intentions to strengthen the ability of the LLMs to reason about complex intentions and user emotions in multi-round interaction scenarios.
 DeepSeek-R1 Guo et al.
@@
 (2025) uses budget forcing to control the number of thinking tokens.
-The upper limit is terminated early by a delimiter, while the lower limit prohibits delimiters and adds ”wait” to guide in-depth reasoning and optimize the quality of the answer.
-SoulChat Chen et al.
-(2023) fine-tuned LLMs by constructing a dataset containing more than 2 million samples of multi-round empathic conversations.
+The upper limit is terminated early by a delimiter, while the lower limit prohibits delimiters and adds "wait" to guide in-depth reasoning and optimize the quality of the answer.
 MoChat Mo et al.
 (2024) constructs multi-round dialogues for spatial localization by using joint grouping spatio-temporal localization.
-4 Realiable Generation
+Cat. Benchmark Year Lang. Num. Type Description
+Instrution Understanding BotChat Duan et al. (2024) 2023 En&Zh 7658 M Multi-round dialogue eval. via simulated data
+MINT Wang et al. (2024) 2023 En 29,307 M Tool use and feedback in multi-turn dialogue
+MT-BENCH-101 Bai et al. (2024) 2024 En 1,388 M Multi-turn dialogue ability
+$\infty${B}ench Zhang et al. (2024c) 2024 En&Zh 100k S Long-context handling
+L-Eval An et al. (2024) 2024 En 2,000 S Evaluation of Long-Context Language Models
+LongICLBench Li et al. (2025) 2025 En 2,100k S Long In-Context Learning
+Intention Reasoning BIPIA Yi et al. (2025) 2023 En 712.5K S Vulnerability to hint injection
+Miko Lu et al. (2024) 2024 En 10k S Multimodal social intent understanding
+CONTRADOC Li et al. (2024a) 2023 En 891 S Self-contradiction in long docs
+CDCONV Zheng et al. (2022) 2022 Zh 12K M Contradiction in Chinese dialogues
+Relibale Dialog Generate Open-LLM-Leaderboard Ye et al. (2024) 2024 En 10K S Uncertainty in generation
+ETHICS Hendrycks et al. (2021) 2020 En 130K S Moral reasoning
+FACTOR Muhlgay et al. (2024) 2023 En 300 S Factuality in generated text
+Table 1: A selection of widely used benchmark datasets for evaluating LLMs. Th“Cat.”: task stage; ’Lang.’: language of the benchmark, a; ’Num.’: data size; ’Type’: S=Single-round, M=Multi-round.
+4 Reliable Dialog Generation
 Despite strong performance, LLMs struggle with output reliability.
@@
 While effective on familiar data, they often produce unstable or incomplete responses to unseen inputs, undermining reliability.
-We analyze LLM response stability and alignment in depth.
 4.1 Response Stability
@@
 If knowledge samples that the model has not seen are used in instruction-tuning, it will inevitably cause the model to give a definite response to unknown inputs, and there is a high probability that an answer will be fabricated.
-This is the over-confidence of the model that causes the model to output unreliable answers, and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Fabricated incorrect information (Figure 5 A.1).
+This is the over-confidence of the model that causes the model to output unreliable answers, and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Fabricated incorrect information (Figure 10 Q.1).
 When the model’s knowledge did not match the input question, it fabricated information that did not match the facts.
-2) Incorrect context output (Figure 5 A.2).
+2) Incorrect context output (Figure 10 Q.2).
 When the model’s knowledge did match the input question, it output incorrect context information.
-To address these issues, researchers have explored uncertainty, which quantifies the credibility and stability of model outputs as follows.
-Figure 5: Common generation issues in LLMs include: (A) Unstable Content Generation (§4.1), where the model fabricates details when it lacks relevant knowledge (A.1) or produces incorrect contextual information despite having relevant knowledge (A.2).
-(B) Misalignment with Human Values (§4.2), where the model generates harmful or offensive content (B.1) or provides responses that conflict with moral and ethical standards (B.2).
+To address these issues, researchers have explored uncertainty, which quantifies the credibility and stability of model outputs.
 Fine-tuning LLMs.
@@
 (2024).
+ConfidenceElicitation Xiong et al.
+(2024) is a new uncertainty measurement tool for large model outputs.
 CalibrateMath Lin et al.
-(2022) assesses uncertainty by requiring models to generate numerical answers with confidence levels, evaluating their reliability.
+assesses uncertainty by requiring models to generate numerical answers with confidence levels, evaluating their reliability.
 4.2 Alignment
@@
 (2021).
-This means that the models may inadvertently produce outputs misaligned with human values as follows: 1) Generation of Toxic Content (Figure 5 B.1).
+This means that the models may inadvertently produce outputs misaligned with human values as follows: 1) Generation of Toxic Content (Figure 11 Q.1).
 LLMs may generate toxic content, such as hate speech or offensive comments, when asked to respond to sensitive topics Luong et al.
@@
 (2024).
-2) Conflicts with Moral/Ethical Standards (Figure 5 B.2).
+2) Conflicts with Moral/Ethical Standards (Figure 11 Q.1).
 LLMs might produce outputs that conflict with moral or ethical standards, such as guiding illegal activities Ramezani and Xu (2023); Abdulhai et al.
 (2024).
-To tackle the previously mentioned concerns regarding unsafe or harmful content produced by LLMs, researchers have focused on various stages:
+To tackle the above concerns regarding unsafe or harmful content produced by LLMs, researchers have focused on various stages:
 Pretraining Data Cleaning and Curation.
@@
 can help identify and remove toxic and biased content.
-Table 1: A selection of widely used benchmark datasets for evaluating LLMs.
-The ’Type’ column categorizes the benchmarks based on task types, while ’Lang.’ indicates the language of the dataset, and ’Citation’ denotes the number of citations each benchmark has received.
-Type Benchmark Year Lang. Description Citation
-[…]
-MT-BENCH-101 Bai et al. (2024) 2024 En Evaluation of multi-turn dialogue capability in LLMs. 39
-LLM Reasoning BIPIA Yi et al. (2023) 2023 En Assesses LLMs vulnerability to hint injection attacks. 60
-Miko Lu et al. (2024) 2024 En Evaluates multimodal LLMs in understanding social intention. 2
-CONTRADOC Li et al. (2024a) 2023 En Detects self-contradictions in long documents. 7
-CDCONV Zheng et al. (2022) 2022 Zh Detects contradictions in Chinese dialogues. 12
-LLM Generation Open-LLM-Leaderboard Ye et al. (2024) 2024 En Assesses LLMs uncertainty in generated content. 19
-ETHICS Dataset Hendrycks et al. (2021) 2020 En Evaluates LLMs on moral reasoning and ethics. 468
-FACTOR Muhlgay et al. (2024) 2023 En Measures factual accuracy in LLM-generated text. 71
 Reinforcement Learning-based Alignment.
-To further align LLMs with human values and societal norms, reinforcement learning methods, such as RLHF and its advanced variants, such as PPO Ouyang et al.
+To further align LLMs with human and societal norms, reinforcement learning methods, such as RLHF and its advanced variants, such as PPO Ouyang et al.
 (2022), DPO Zeng et al.
@@
 In-context Alignment.
-In-context alignment leverages the ability of LLMs to adapt their responses based on a few examples provided in the prompt.
+It leverages the ability of LLMs to adapt their responses based on a few examples provided in the prompt.
 Lin et al.
@@
 (2024a) explored the effectiveness of different components of In-context alignment, and found that examples within the context are crucial for enhancing alignment capabilities.
+5 Case Analysis
+To demonstrate the challenges of existing LLMs in above stages, we provide cases analysis in Appendix B, and make statistical comparisons in Table 2.
+The results show that models with reasoning capabilities (such as GPT-o3*, Deepseek-R1*) perform better in command understanding tasks such as long text understanding and multi-round dialogue, but have not completely solved all problems.
+In the intention reasoning task, except for "Inconsistent information reasoning", the models generally performed poorly, and only a few models partially passed the case.
+In terms of reliable dialogue generation, all models performed poorly in response stability, and only GPT-o3 and Deepseek-R1 performed well in the alignment test.
+Overall, while reasoning abilities contribute to improved task performance, significant limitations remain across all three stages.
 6 Benchmark
-This section covers benchmarks for LLMs in instruction understanding, reasoning, and generation (Table 1).
+This section covers benchmarks for LLMs in above three stage (Table 1), more details see Appendix A.
 6.1 Benchmarking Instruction Understanding
 Instruction understanding in LLMs involves extracting key information, maintaining coherence, and adapting to dynamic conversation changes, especially in long or multi-round dialogues.
+LLM capabilities in multi-turn dialogue and long-context processing have been explored through various benchmarks.
+BotChat Duan et al.
+(2024) evaluates dialogue generation, showing GPT-4’s strengths but noting instruction compliance and length limitations in other models.
+MINT Wang et al.
+(2024) highlights limited progress in tool use and feedback for complex tasks.
+MT-Bench-101 Bai et al.
+(2024) identifies challenges in enhancing long-term interaction skills.
+For long-context tasks, performance drops in ultra-long texts (Zhang et al.
+Zhang et al.
+(2024c)), L-Eval An et al.
+(2024) emphasizes length-instruction-enhanced metrics, and LongICLBench Li et al.
+(2025) reveals difficulties in reasoning across multiple pieces of information.
+Stage Challenge GPT-4o GPT-o3* Qwen3 Qwen3* Deepseek-v3 Deepseek-R1*
+Instruction Understanding Long-Text Comprehension C A C A C A
+Multi-Turn Conversation Handling C A C C C A
+Intention Reasoning Inconsistent Instruction Reasoning C C C C B C
+Misinformation Reasoning B A C B B A
+Fuzzy Language Interpretation B B C C C C
+Intention Clarification Failure C B C C C C
+Reliable Dialog Generation Response Stability C C C C C C
+Alignment C A C C C A
+Table 2: Results of using different LLMs on the challenge cases; * denotes models with reasoning abilities.
+’A’ indicates the model’s accuracy is greater than 75%, ’B’ is between 50% and 75%, and ’C’ is below 50%.
+6.2 Benchmarking LLM Reasoning
+LLM intention reasoning involves inferring user intentions by interpreting both explicit and implicit language cues.
+The existing benchmarks comprehensively evaluate the multifaceted reasoning capabilities of LLMs, encompassing vulnerability to indirect hint injection attacks (BIPIA Yi et al.
+(2025)), advantages in multimodal intention understanding (Miko Lu et al.
+(2024)), analysis of self-contradictions in long documents (CONTRADOC Li et al.
+(2024a)), and contradiction detection in Chinese dialogues (CDCONV Zheng et al.
+(2022)).
+Collectively, these benchmarks highlight both the challenges and advancements of LLMs in complex reasoning.
+6.3 Benchmarking LLM Generation
+LLM Generation assesses a model’s ability to understand user instructions, avoid fabricating false information, and generate accurate, contextually appropriate responses.
+Open-LLM-Leaderboard Ye et al.
+(2024), ETHICS Hendrycks et al.
+(2021) and FACTOR Muhlgay et al.
+(2024) all focus on the reliability evaluation of content generated by large models.
+Open-LLM-Leaderboard finds that large-scale models have higher uncertainty, and fine-tuned models have higher accuracy but greater uncertainty.
+ETHICS focuses on the ethical value alignment of generated content.
+FACTOR evaluates factuality through scalable methods to ensure that diverse and rare facts are covered.
+7 Future Directions
+This section summarizes ongoing challenges in above stages with LLMs and outlines potential future research directions.
+Automated Annotation Framework.
+Although LLMs excel in general-domain tasks, they often produce hallucinated or incomplete content in specialized fields due to limited domain-specific training data.
+While contextual learning and instruction fine-tuning methods have been explored to address this issue, manual data annotation remains labor-intensive and prone to quality inconsistencies.
+An automated annotation framework could streamline data labeling, enhancing model performance in specialized fields by ensuring higher quality and scalability of domain-specific training datasets.
+GraphRAG.
+LLMs have shown impressive language generation capabilities through pre-training on large datasets, but their reliance on static data often results in inaccurate or fictional content, particularly in domain-specific tasks.
+The Graph-enhanced generation approach aims to tackle this by leveraging KGs and GNNs for precise knowledge retrieval.
+Despite its advantages, GraphRAG faces challenges in capturing structural information during graph reasoning tasks and struggles with multi-hop retrieval accuracy and conflict resolution between external and internal knowledge.
+Future work should focus on refining retrieval strategies and improving the stability and accuracy of GraphRAG in complex tasks.
+Balancing Safety and Performance.
+Although advancements in alignment techniques have improved factual accuracy and safety, they often come at the cost of the model’s creativity and fluency.
+Striking a balance between safety and performance is crucial.
+Future research should explore new alignment methods that ensure both the safety and usability of LLMs, optimizing the trade-off between generating reliable, safe content and maintaining the model’s creative and contextual capabilities.
+8 Conclusion
+This paper analyzes LLMs’ performance in processing user instructions.
+Despite progress in natural language understanding, LLMs struggle with complex, inconsistent instructions, often resulting in biases, errors, and hallucinations.
+Improvements through prompt engineering, model expansion, and RLHF et al.
+have not fully addressed LLMs’ limitations in reasoning and comprehension, limiting their real-world applicability.
+We identify three challenges: instruction understanding, intention reasoning and reliable dialog generation.
+Future research should focus on enhancing reasoning for complex instructions and aligning outputs with user intention to improve LLMs’ reliability.
+Limitations
+We have made significant efforts to ensure the quality of this study, but certain limitations may still exist.
+First, due to constraints on the length of the discussion, this article primarily focuses on analyzing examples of human interaction with large models, which inevitably limits the depth of methodological details we can explore.
+Second, while our meta-analysis draws extensively from prominent academic platforms such as the ACL conference series, ICLR, ICML, and the arXiv preprint repository, it is possible that some valuable insights from other sources have not been included.
+It is worth noting that several open scientific questions in this field remain unresolved, and the academic community has yet to reach a consensus on these issues.
+To address these limitations, we plan to establish a long-term tracking mechanism to monitor new developments in the field.
+This will allow us to incorporate emerging theoretical advancements and dynamically refine or expand upon the perspectives presented in this study.
+Appendix A Benchmark Details
+A.1 Instruction Understanding Benchmark
+BotChat Duan et al.
+(2024) specifically addresses the evaluation of LLMs’ ability to emulate human-like, multi-turn conversations using an LLM-centric approach.
+This benchmark assesses utterance generation, evaluation protocols, and judgment.
+Findings suggest that models like GPT-4 demonstrate exceptional performance as both generators and judges, producing human-indistinguishable dialogues and showing high alignment with human evaluation standards.
+Conversely, other LLMs face challenges in generating quality multi-turn dialogues due to issues like inadequate instruction-following and a tendency towards prolixity, particularly in generating extensive dialogues.
+MINT Wang et al.
+(2024) introduces a benchmark that evaluates LLMs’ capacity to solve complex tasks through multi-turn interactions, emphasizing the use of external tools and leveraging natural language feedback.
+MINT provides a reproducible evaluation framework where LLMs can access tools and receive simulated user feedback.
+Analysis using MINT reveals that LLMs generally benefit from tools and language feedback.
+Interestingly, studies with MINT suggest that techniques like supervised instruction-finetuning (SIFT) and reinforcement learning from human feedback (RLHF) might not consistently improve multi-turn capabilities, indicating the need for further research in training methodologies for conversational settings.
+MT-BENCH-101 Bai et al.
+(2024) offers a structured approach with its three-layer hierarchical capability taxonomy and a dataset of 1,388 dialogue pairs across 13 tasks.
+This allows for a fine-grained evaluation of LLMs’ multi-turn dialogue skills.
+The findings from studies utilizing MT-BENCH-101 indicate that commonly employed techniques yield limited improvements in multi-turn performance, suggesting a need for more effective methods to enhance conversational abilities over extended interactions.
 Zhang et al.
 Zhang et al.
-(2024c) provides a benchmark consisting of 12 tasks, with an average data length exceeding 100K tokens, designed to evaluate LLMs’ ability to handle long contexts.
-The results show that the performance of LLMs significantly declines when dealing with long contexts, suggesting further improvements.
-MT-BENCH-101 Bai et al.
-(2024) constructs a three-layer hierarchical capability classification and 1,388 dialogue pairs across 13 tasks, providing a comprehensive evaluation of LLMs’ fine-grained capabilities in multi-turn dialogue.
-Experimental results show that commonly used techniques do not significantly enhance multi-turn capabilities.
-5.2 Benchmarking LLM Reasoning
-LLM Reasoning involves inferring user intentions by interpreting both explicit and implicit language cues.
+(2024c) focus on the challenge of long-context processing, introducing a benchmark of 12 tasks, each averaging over 100K tokens, to assess LLMs’ long-context processing abilities.
+Results indicate a marked performance drop with longer contexts, highlighting the need for further improvement.
+L-Eval An et al.
+(2024) contributes to standardizing the evaluation of Long-Context Language Models (LCLMs) by providing a new evaluation suite with diverse tasks, document lengths (3k to 200k tokens), and a large number of human-labeled query-response pairs.
+It also investigates the effectiveness of evaluation metrics, suggesting that Length-instruction-enhanced (LIE) evaluation and LLM judges correlate better with human judgments.
+Comprehensive studies using L-Eval offer insights into the performance of both commercial and open-source LCLMs.
+LongICLBench Li et al.
+(2025) focuses on long in-context learning in extreme-label classification, with input lengths ranging from 2K to 50K tokens and a large number of classes.
+Evaluation on this benchmark showed that while LLMs perform well on less challenging tasks, they struggle with more complex ones involving a large label space and longer contexts, revealing a bias towards later parts of the input and difficulty in reasoning over multiple pieces of information.
+A.2 LLM Intention Reasoning Benchmark
 BIPIA Yi et al.
@@
 The study shows that the Hierarchical method consistently outperforms others in detecting contradictions, highlighting the importance of accurate contextual modeling in dialogue understanding.
-5.3 Benchmarking LLM Generation
-LLM Generation assesses a model’s ability to understand user instructions, avoid fabricating false information, and generate accurate, contextually appropriate responses.
+A.3 LLM Reliable Generation Benchmark
 Open-LLM-Leaderboard Ye et al.
@@
 Testing models such as GPT-2 and GPT-Neo show that, while benchmark scores correlate with perplexity, they better reflect factuality in open-ended generation, especially when retrieval augmentation is applied.
-6 Future Directions
-This section summarizes ongoing challenges in instruction understanding, reasoning, and reliable generation with LLMs and outlines potential future research directions.
-Automated Annotation Framework.
-Although LLMs excel in general-domain tasks, they often produce hallucinated or incomplete content in specialized fields due to limited domain-specific training data.
-While contextual learning and instruction fine-tuning methods have been explored to address this issue, manual data annotation remains labor-intensive and prone to quality inconsistencies.
-An automated annotation framework could streamline data labeling, enhancing model performance in specialized fields by ensuring higher quality and scalability of domain-specific training datasets.
-GraphRAG.
-LLMs have shown impressive language generation capabilities through pre-training on large datasets, but their reliance on static data often results in inaccurate or fictional content, particularly in domain-specific tasks.
-The Graph-enhanced generation approach aims to tackle this by leveraging KGs and GNNs for precise knowledge retrieval.
-Despite its advantages, GraphRAG faces challenges in capturing structural information during graph reasoning tasks and struggles with multi-hop retrieval accuracy and conflict resolution between external and internal knowledge.
-Future work should focus on refining retrieval strategies and improving the stability and accuracy of GraphRAG in complex tasks.
-Quantifying Uncertainty in LLMs.
-As LLMs like LLama and ChatGPT have revolutionized content generation, ensuring the reliability of their outputs remains a critical challenge.
-Uncertainty quantification is a promising approach to address this, enabling models to provide a confidence assessment alongside their responses.
-Evidence learning Sensoy et al.
-(2018); Amini et al.
-(2020), an emerging method in uncertainty representation, offers a reliable approach for quantifying uncertainty directly from data.
-However, its application to LLMs is computationally intensive, and most existing work focuses on small-scale models.
-Future research should aim to optimize uncertainty quantification for large-scale models efficiently.
-Balancing Safety and Performance.
-Although advancements in alignment techniques have improved factual accuracy and safety, they often come at the cost of the model’s creativity and fluency.
-Striking a balance between safety and performance is crucial.
-Future research should explore new alignment methods that ensure both the safety and usability of LLMs, optimizing the trade-off between generating reliable, safe content and maintaining the model’s creative and contextual capabilities.
-7 Conclusion
-This paper analyzes LLMs’ performance in processing user instructions.
-Despite progress in natural language understanding, LLMs struggle with complex, inconsistent instructions, often resulting in biases, errors, and hallucinations.
-Improvements through prompt engineering, knowledge retrieval, model expansion, fine-tuning, and RLHF have not fully addressed LLMs’ limitations in reasoning and comprehension, limiting their real-world applicability.
-We identify three key bottlenecks: instruction comprehension, reasoning ability, and reliable content generation.
-Future research should focus on enhancing reasoning for complex instructions and aligning outputs with user intent, to improve LLMs’ adaptability and reliability in real-world tasks.
+Appendix B Challenges Cases in Human-LLMs Aligement
+To fully understand the limitations of large language models (LLMs) in practical applications, it is particularly important to analyze their performance in a variety of complex scenarios.
+We collected human user instructions from real scenarios that the large model needs to interact with on a total of eight challenges that the summarized existing large language models (LLMs) face in the three phases of instruction understanding, intention reasoning, and reliable dialog generation, and cleaned and filtered the instructions through manual screening, diversity sampling, and difficulty filtering, and finally, 50 human instructions were used for each challenge.
+The statistical results of the test are shown in Table.2.
+In this section, we will present a series of representative case studies and conduct an in-depth analysis of the current mainstream LLMs (including GPT-4o, GPT-o3, Qwen3, DeepSeek-V3 and DeepSeek-R1).
+Besides, the blue smiley face indicates that the model is capable of providing accurate responses or identifying inconsistencies in the user’s input instructions.
+In contrast, the red crying face signifies that the LLM failed to recognize contradictions in the user’s instructions and produced incorrect responses.
+Through these cases, we systematically reveal the typical failure modes of each model in different aspects and their deep-seated causes, providing important references and directions for targeted optimization in future research.
+B.1 Case of Instruction Understanding
+Although large language models (LLMs) perform well in short text and single-round dialogue scenarios, their ability to understand and execute complex instructions in long contexts and multi-round dialogues still faces many severe challenges.
+Figure 4 shows that long texts usually contain a lot of irrelevant or redundant information, which causes the core content directly related to the task to be submerged.
+The model is prone to omission or confusion when extracting key information, and even hallucinations, thereby generating content that is irrelevant to the user’s instructions.
+Figure 5 shows that when dealing with long-distance information associations that need to span multiple paragraphs or dialogue rounds, the model often has difficulty tracking and integrating the logical relationship between contexts, thereby losing important clues.
+In multi-round dialogue scenarios, the model is not only prone to gradually accumulate errors due to inaccurate understanding of the previous text, but may also fail to correctly judge the relevance of each round of dialogue, causing the answer to deviate from the user’s real needs.
+As shown in Table 2, Figure 4, and Figure 5, in the instruction understanding stage, the reasoning models generally performed well on the "Long-Text Comprehension" task, while the non-reasoning models all struggled, and for “Multi-Turn Conversation Handling”, in addition to the non-reasoning models, Qwen3, which provides reasoning capability, also exhibited suboptimal performance.
+The above results show that reasoning can effectively improve the ability of LLMs in instruction understanding, but cannot completely solve the instruction understanding problem.
+B.2 Case of Intention Reasoning
+Since there are often spelling errors, factual contradictions and semantic ambiguities in user instructions, LLM faces many challenges in understanding the user’s true intentions.
+Figure 6 shows that the model often relies too much on the user’s input and ignores the obvious errors in the input.
+It tends to generate answers directly instead of identifying and correcting these errors first;
+Figure 7 shows that when knowledge updates are not synchronized or input data is maliciously tampered with, the model is more likely to output information that is inconsistent with actual needs or contaminated.
+However, the reasoning model can identify the errors, indicating that the model’s reasoning ability is crucial in identifying the user’s input intention.
+Figure 8 shows that when faced with ambiguous or uncertain expressions, large models usually tend to customize questions or give default explanations based on their own training preferences, rather than actively asking users for more context.The GPT series performs better than models such as DeepSeek in such tasks, indicating that even models with strong reasoning capabilities may have difficulties in dealing with modal expressions such as sarcasm and irony.
+This reflects the limitations of AI in understanding complex human language expressions, as well as differences in the coverage of such language phenomena in the training data and the depth of the model’s understanding of the social and cultural context.
+Figure 9 shows that for implicit intentions such as sarcasm, metaphors, and emotions, the model often only focuses on the literal meaning and has difficulty grasping the deep emotions or context, thereby outputting incorrect analysis results.
+As shown in Table 2, Figure 6 to Figure 9, in the intention reasoning stage, all models encountered difficulties and performed poorly in “Inconsistent Instruction Reasoning”, “Fuzzy Language Interpretation” and “Intention Clarification Failure”; in “Misinformation Reasoning”, only GPT-o3 and Deepseek-R1 achieved good performance, while the other models underperformed.
+This suggests that all models have significant challenges in reasoning about intentions.
+B.3 Case of Reliable Dialog Generation
+Current large language models exhibit both deterministic response preferences and ambiguous knowledge boundaries during generation.
+When queries fall within the model’s knowledge coverage, the outputs are generally reliable; however, in open-domain or previously unseen scenarios, the quality of generated content fluctuates significantly.
+Figure 10 illustrates the instability of LLM-generated content, which primarily manifests in two typical issues: First, when the model lacks relevant knowledge or information regarding the input (Q.1), it tends to fabricate details, producing content inconsistent with facts and thereby substantially compromising the accuracy and reliability of its outputs.
+Second, even when the model possesses relevant knowledge (Q.2), it may still misinterpret or improperly integrate contextual information, resulting in outputs that are contextually inappropriate or logically flawed.
+These issues not only undermine user experience but also limit the applicability of LLMs in high-stakes scenarios.
+Figure 11 demonstrates the misalignment between model-generated content and human values, mainly reflected in the model’s potential to violate widely accepted moral and ethical standards.
+On one hand, when handling sensitive or controversial topics, the model may generate harmful, offensive, or discriminatory statements, negatively impacting users.
+On the other hand, it may also produce responses that encourage illegal activities or contravene social ethics, introducing potential legal and societal risks and imposing higher requirements on the safety and trustworthiness of LLMs.
+As shown in Table 2, Figure 10 and Figure 11, in the reliable dialog generation stage, all models failed to maintain “Response Stability”; in “Alignment” test, only GPT-o3 and Deepseek-R1 performed well, while the rest of the models failed.
+The results show that all models have clear shortcomings in “Response Stability”, and the reasoning models have improved in “Alignment”.
+Overall, reasoning capabilities can improve the performance of llms for interacting with real human user commands, however, all models still face significant challenges and remain underpowered in real-world and human interaction scenarios.
+Figure 4: Case of Remote Information Failure (§2.1), where the model forgets relevant information over long distances in long context.
+Figure 5: Case of Incorrect Relevance Judgment2.2), such as the model incorrectly associates wrong content from the previous turn.
+Figure 6: Case of Inconsistent Instruction Reasoning (§3.1), where LLMs fail to detect conflicting inputs (Q.1) or overlooking logical inconsistencies (Q.2)
+Figure 7: Case of Misinformation Reasoning3.2), caused by temporal misalignment leading to failure responses (Q.1) or data contamination resulting in misleading outputs (Q.2).
+Figure 8: Case of Fuzzy Language Interpretation3.3), where the model relies on biases for fuzzy queries (C.1) or defaults to a response without seeking clarification (C.2)
+Figure 9: Case of Intention Clarification Failure3.4) , where it misinterprets sarcasm (Q.1) or ignores prior emotional context (Q.2)
+Figure 10: Case of Unstable Content Generation (§4.1), where the model fabricates details when it lacks relevant knowledge (Q.1) or produces incorrect contextual information despite having relevant knowledge (Q.2)
+Figure 11: Case of Misalignment with Human Values (§4.2), where the model generates harmful or offensive content (Q.1)

### v3

@@
-A Survey of Progress and Challenges in Aligning LLMs with Human Intentions
+A Survey of LLM Alignment: Instruction Understanding, Intention Reasoning, and Reliable Generation
 Abstract
-Large language models (LLMs) have demonstrated exceptional capabilities in understanding and generation.
-However, when interacting with human instructions in real-world scenarios, LLMs still face significant challenges, particularly in accurately capturing and comprehending human instructions and intentions.
+Large language models have demonstrated exceptional capabilities in understanding and generation.
+However, in real-world scenarios, users’ natural language expressions are often inherently fuzzy, ambiguous, and uncertain, leading to challenges such as vagueness, polysemy, and contextual ambiguity.
 This paper focuses on three challenges in LLM-based text generation tasks: instruction understanding, intention reasoning, and reliable dialog generation.
@@
 Finally, we explore potential directions for future research to enhance the reliability and adaptability of LLMs in real-world applications.
-A Survey of Progress and Challenges in Aligning LLMs
-with Human Intentions
-Zongyu Chang1thanks: These authors contributed equally to this work., Feihong Lu2*, Ziqin Zhu2, Qian Li1, Cheng Ji2, Zhuo Chen1, Hao Peng2, Yang Liu5, Ruifeng Xu3, Yangqiu Song4, Jianxin Li2, Shangguang Wang1 1Beijing University of Posts and Telecommunications, 2Beihang University, 3Harbin Institute of Technology, Shenzhen, 4Hong Kong University of Science and Technology, 5Chinese Academy of Sciences {changzongyu, li.qian, sgwang}@bupt.edu.cn, {lufeihong, zhuziqin, jicheng, lijx}@buaa.edu.cn, xuruifeng@hit.edu.cn, yqsong@cse.ust.hk
+keywords:
+Instruction Understanding, Intention Reasoning, Reliable Dialog Generation
+11footnotetext: * These authors contributed equally to this work.
+[…]
+[…]
 1 Introduction
@@
 These models, built upon massive amounts of data and extensive computing resources, have shown impressive capabilities in understanding and generating human language.
-Recent advancements in LLMs, including the use of scaling laws Kaplan et al.
-(2020), supervised fine-tuning (SFT) Wu et al.
-(2021), and reinforcement learning with human feedback (RLHF) Ouyang et al.
-(2022), have propelled these models to new heights.
-Researchers have explored innovative strategies like chain-of-thought reasoning (COT) Wei et al.
-(2022), aiming to enhance their performance in processing and generating accurate responses.
-However, they still struggle with more complex interactions, especially when the input data is ambiguous, incomplete, or inconsistent.
-Despite improvements, issues such as content hallucination Li et al.
-(2023) and logical misinterpretations remain prevalent.
+Recent advancements in LLMs, including the use of scaling laws [1], supervised fine-tuning (SFT) [2], and reinforcement learning with human feedback (RLHF) [3], have propelled these models to new heights.
+Researchers have explored innovative strategies like chain-of-thought reasoning (COT) [4], aiming to enhance their performance in processing and generating accurate responses.
+However, they still have struggled when interacting with human instructions in real-world scenarios due to the fuzzy, uncertain, and complex nature of human instructions, especially when the input data is fuzzy, incomplete, or inconsistent.
+Despite improvements, issues such as content hallucination [5] and logical misinterpretations remain prevalent.
 Consequently, while LLMs show promise, they are far from flawless and require further refinement to address the challenges posed by more unpredictable and complex human instructions as follows.
@@
 Figure 2: Unlike previous surveys on LLMs, we do not consider the alignment between LLMs and humans as an isolated process,instead, we view it as a continuous and dynamic information processing process consisting of instruction understanding, intention reasoning, and reliable dialogue generation.
-I.
-Challenge of Instruction Understanding.
+1.1 Challenge of Instruction Understanding
 One of the most pressing challenges that LLMs face is instruction understanding as Figure 1(a) and Figure (2I), particularly when the user input involves complex or multi-step instructions.
@@
 LLMs often fail to grasp subtle nuances or interpret implicit meanings embedded within the text, which leads to inaccurate or incomplete responses.
-Existing approaches to instruction understanding have introduced techniques like optimizing the model’s parsing abilities Teng et al.
-(2024), and context-aware optimization Sun et al.
-(2024).
+Existing approaches to instruction understanding have introduced techniques like optimizing the model’s parsing abilities [6], and context-aware optimization [7].
 While these methods show promise, they often fall short when addressing the complexities and ambiguities present in instructions.
-II.
-Challenge of Intention Reasoning.
+1.2 Challenge of Intention Reasoning
 Another critical area is intention reasoning as illustrated in Figure 1(b) and Figure 2 (II), where they struggle to align the generated responses with the user’s underlying intention.
@@
 However, these methods often introduce new challenges related to bias and the inability to fully resolve conflicts in user input, further complicating the alignment between generated content and user expectations.
-III.
-Challenge of Reliable Dialog Generation.
+1.3 Challenge of Reliable Dialog Generation.
 The final major challenge is the reliable dialog generation, which pertains to the accuracy, ethical considerations, and stability of the content they produce, such as Figure 1(c) and Figure 2 (III.
@@
 {forest}
-forked edges, for tree= grow=east, reversed=true, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, anchor=center , [ Gap Between LLMs and Human Intention, ver, color=hidden-draw, fill=mygray!30, text=black [ Instruction
-Understanding (§2), color=myblue!100, fill=myblue!80, text width=6em, text=black [ Long-text
-Comprehension (§2.1), color=bluelight1!60, fill=bluelight1!60, text width=7em, text=black [ Information
-Focusing, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ 1) Attention Sparsification Beltagy et al. (2020),
-2) Attention Optimization Chen et al. (2024),
-3) Position independent Training He et al. (2024a) , understanding_leaf, text width=26.5em ] ] [ Multipath
-Optimization, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ 1) SFT and RL Zhang et al. (2024b)
-2) Retrieval-Augmented Li et al. (2024b)
-3) Recurrent Sequence Optimization Gu and Dao (2023) , understanding_leaf, text width=26.5em ] ] ] [ Multi-turn Conversation
-Handling (§2.2), color=bluelight1!60, fill=bluelight1!60, text width=7em, text=black [ Soft
-Fine-tuning, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ Multi-turn SFT Teng et al. (2024), Parrot Sun et al. (2024) , understanding_leaf, text width=26.5em ] ] [ Reforcement
-Learning, color=bluelight1!40, fill=bluelight1!40, text width=4em, text=black [ ArCHer Zhou et al. (2024) , understanding_leaf, text width=26.5em ] ] ] ] [ Intention
-Reasoning (§3), color=mygreen!80, fill=mygreen!80, text width=6em, text=black [ Inconsistent Instruction
-Resolution (§3.1), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Knowledge
-Updating, color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ SituatedQA Zhang and Choi (2021), CDConv Zheng et al. (2022) , reasoing_leaf, text width=26.5em ] ] [ Confidence
-Calibration, color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ CD2 Jin et al. (2024), MacNoise Hong et al. (2024) , reasoing_leaf, text width=26.5em ] ] ] [ Misinformation
-Reasoning (§3.2), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Targeted
-Fine-tuning, color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ CKL Jang et al. (2022), StreamingQA Liska et al. (2022), BIPIA Yi et al. (2025),
-KC-LLMs Wang et al. (2023), CAR Weller et al. (2024) , reasoing_leaf, text width=26.5em ] ] ] [ Fuzzy Language
-Interpretation (§3.3), color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Clue
-Engineering , color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ Folkscope Yu et al.
-(2023), Miko Lu et al.
-(2024), ATC Deng et al.
-(2023) , reasoing_leaf, text width=26.5em ] ] ] [ Intention
-Clarification
-Failure (§3.4) , color=mygreen!60, fill=mygreen!60, text width=7em, text=black [ Deep
-Reasoning , color=mygreen!40, fill=mygreen!40, text width=4em, text=black [ DeepSeek-R1 Guo et al. (2025), S1 Muennighoff et al. (2025),
-SoulChat Chen et al. (2023), MoChat Mo et al. (2024) , reasoing_leaf, text width=26.5em ] ] ] ] [ Reliable
-Generation (§4), color=lightkhaki!80, fill=lightkhaki!80, text width=6em, text=black [ Response Stability (§4.1), color=lightkhaki!60, fill=lightkhaki!60, text width=7em, text=black [ Fine-tuning
-LLMs, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ EDL Sensoy et al. (2018), DER Amini et al. (2020),
-ConformalFactuality Mohri and Hashimoto (2024), LUQ Zhang et al.
-(2024a) , reliable_leaf, text width=26.5em ] ] [ External Tools, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ SUE Liu et al.
-(2024), CalibrateMath Lin et al.
-, reliable_leaf, text width=26.5em ] ] ] [ Alignment (§4.2), color=lightkhaki!60, fill=lightkhaki!60, text width=7em, text=black [ Data Cleaning
-and Curation, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ Stochastic Parrots Bender et al.
-(2021) , DeepSoftDebias Rakshit et al.
-, reliable_leaf, text width=26.5em ] ] [ RL-based
-Alignment , color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ PPO Ouyang et al. (2022), DPO Zeng et al. (2024), GRPO Shao et al. (2024),
-RLAIF Lee et al. (2024) , reliable_leaf, text width=26.5em ] ] [ In-context
-Alignment, color=lightkhaki!40, fill=lightkhaki!40, text width=4em, text=black [ URIAL Lin et al. (2024), ICA Huang et al. (2024a) , reliable_leaf, text width=26.5em ] ] ] ] ]
 Figure 3: Challenges and existing solutions between LLMs and Human Intentions.
-Present Survey.
+Present Survey
 Facing these challenges, there is an increasing need for focused research on LLMs and their interaction with human instructions and intentions.
 This paper systematically analyzes LLMs’ performance in processing human instructions, highlighting three key areas: user instruction understanding, intention comprehension and reasoning, and reliable dialog generation.
-While existing review papers address model training, fine-tuning, and specific aspects of LLMs’ capabilities Lou et al.
-(2024); Plaat et al.
-(2024); Huang et al.
-(2024b), our focus is on the LLMs’ ability to understand and reason about user intentions.
+While existing review papers address model training, fine-tuning, and specific aspects of LLMs’ capabilities [92, 93, 94], our focus is on the LLMs’ ability to understand and reason about user intentions.
 Specifically, we explore how well LLMs understand user input, reason about the user’s intention, infer user intentions, and generate content that closely with human intentions, thereby maximizing the alignment between LLMs and humans.
-Comparison with Previous Surveys.
+Comparison with Previous Surveys
 While the gap between human intention and LLMs is a core challenge in generative AI, many studies focus on specific aspects of the issue, lacking a comprehensive overview.
@@
 Lou et al.
-Lou et al.
-(2024) primarily address instruction following challenges in LLMs without delving into the reasoning capabilities for complex user instructions.
-Gao et al.
-analyze the four stages of human-machine LLM interaction (planning, facilitation, iteration, and testing) but overlook LLM’s understanding of user instructions.
+[92] primarily address instruction following challenges in LLMs without delving into the reasoning capabilities for complex user instructions.
 Xu et al.
-Xu et al.
-examine the impact of various memory conflicts on LLM-generated content credibility and performance, yet do not consider reasoning or intention comprehension.
+[95] examine the impact of various memory conflicts on LLM-generated content credibility and performance, yet do not consider reasoning or intention comprehension.
 Plaat et al.
-Plaat et al.
-(2024) focus on LLM Reasoning for basic mathematical problems, without exploring its applicability to broader fields.
+[93] focus on LLM Reasoning for basic mathematical problems, without exploring its applicability to broader fields.
 Shorinwa et al.
-Shorinwa et al.
-(2024) provide an initial analysis of LLMs uncertainty quantification, but exclude user input instructions.
+[96] provide an initial analysis of LLMs uncertainty quantification, but exclude user input instructions.
 In contrast, our survey offers a more comprehensive perspective, as shown in Figure 2 and Figure 3, with a unique classification and systematic analysis of instruction processing, while addressing current solutions to key challenges.
-Survey Organization.
+Survey Organization
 As in Figure 3, we begin by exploring the capability of user instruction understanding (§2).
@@
 LLMs may forget prior information, be influenced by irrelevant data, and overlook key inputs.
+Figure 4: Case of Remote Information Failure (§2.1), where the model forgets relevant information over long distances in long context.
 2.1 Long-Text Comprehension
@@
 This paper classifies the existing solutions into the following two categories:
-Information Focusing.
-Improving LLM’s ability to focus on important information in long texts involves several methods: 1) Sparsifying attention to concentrate on critical information Beltagy et al.
-(2020).
-2) Optimizing attention to minimize redundancy and emphasize core content Chen et al.
-(2024).
-3) Training with location-independent tasks to enhance the ability to search and react to relevant information in long contexts He et al.
-(2024a).
-Multipath Optimization.
-Various methods can enhance LLMs on long-context tasks: 1) Pre-training with extended context windows and reinforcement learning for fine-tuning to optimize long-context understanding Zhang et al.
-(2024b).
-2) Combining retrieval-based models with generative models on long-context tasks Li et al.
-(2024b).
-3) Leveraging cyclic sequence models’ linear scaling property for better inference efficiency Gu and Dao (2023).
+Information Focusing
+Improving LLM’s ability to focus on important information in long texts involves several methods: 1) Sparsifying attention to concentrate on critical information [8].
+2) Dynamically adjusting attention based on the current task [9].
+3) Optimizing attention to minimize redundancy and emphasize core content [10].
+4) Training with location-independent tasks to enhance the ability to search and react to relevant information in long contexts [11].
+Multipath Optimization
+Various methods can enhance LLMs on long-context tasks: 1) Pre-training with extended context windows and reinforcement learning for fine-tuning to optimize long-context understanding [12, 13, 14, 15].
+2) Combining retrieval-based models with generative models on long-context tasks [16, 17].
+3) Leveraging cyclic sequence models’ linear scaling property for better inference efficiency [18].
+4) Using external memory to store and retrieve long-context information [19].
+5) Mimicking brain memory hierarchies to improve long-context processing efficiency [20].
+Figure 5: Case of Incorrect Relevance Judgment2.2), such as the model incorrectly associates wrong content from the previous turn.
 2.2 Multi-Turn Conversation Handling
@@
 However, due to the characteristics of real-world conversations, such as the constant changes in user intentions and long-distance dependencies, LLMs still faces significant challenges in achieving coordinated multi-round interactions with humans.This paper categorizes the challenges faced by existing LLMs when understanding multi-turn conversations into three categories, as follows: 1) Capability Weakening.
-Current supervised instruction fine-tuning (SIFT) and RLHF may even impair multi-turn capabilitiesWang et al.
-(2024), with models struggling on complex reasoning tasks that span multiple rounds, such as those requiring evidence collection and conclusions Banatt et al.
-(2024).
-Additionally, multi-turn dialogs increase the vulnerability of LLMs to adversarial attacks, where malicious users can mask harmful intentions across multiple rounds, leading to the generation of misleading or harmful content Agarwal et al.
-(2024).
+Current supervised instruction fine-tuning (SIFT) and RLHF may even impair multi-turn capabilities[97], with models struggling on complex reasoning tasks that span multiple rounds, such as those requiring evidence collection and conclusions [98].
+Additionally, multi-turn dialogs increase the vulnerability of LLMs to adversarial attacks, where malicious users can mask harmful intentions across multiple rounds, leading to the generation of misleading or harmful content [99].
 2) Error Propagation.
-Instruction comprehension errors accumulate across rounds, leading to an escalating failure rate in subsequent responses He et al.
-(2024b), which may snowball into larger issues such as biased or incorrect outputs Fan et al.
-(2024).
-3) Incorrect Relevance Judgment (Figure 5 Q.1) LLMs often struggle to identify relevant content in multi-turn dialogs, failing to properly link content from previous rounds or to discern ellipsis and implicit meaning inherent in user commands Sun et al.
-(2024).
-To solve above challenges, this paper categorizes existing solutions into two types: supervised fine-tuning methods using multi-turn dialogue data, enhanced by techniques like optimized instruction parsing Teng et al.
-(2024) and context-aware preference strategies Sun et al.
-(2024); and reinforcement learning methods tailored for multi-turn dialogue, with improvements such as hierarchical reinforcement learning Zhou et al.
-(2024).
+Instruction comprehension errors accumulate across rounds, leading to an escalating failure rate in subsequent responses [100], which may snowball into larger issues such as biased or incorrect outputs [101].
+3) Incorrect Relevance Judgment (Figure 5 Q.1) LLMs often struggle to identify relevant content in multi-turn dialogs, failing to properly link content from previous rounds or to discern ellipsis and implicit meaning inherent in user commands [7].
+To solve above challenges, this paper categorizes existing solutions into two types: supervised fine-tuning methods using multi-turn dialogue data, enhanced by techniques like optimized instruction parsing [6] and context-aware preference strategies [7]; and reinforcement learning methods tailored for multi-turn dialogue, with improvements such as hierarchical reinforcement learning [28].
 3 Intention Reasoning
@@
 We explore causes and solutions for intention errors, focusing on inconsistent instructions, misinformation, fuzzy language, and intention clarification.
+Figure 6: Case of Inconsistent Instruction Reasoning (§3.1), where LLMs fail to detect conflicting inputs (Q.1) or overlooking logical inconsistencies (Q.2)
 3.1 Inconsistent Instruction Reasoning
 In natural language communication, humans easily identify inconsistencies using context and prior knowledge, whereas LLMs struggle, often accept contradictory inputs, and generate unreliable answers.
-This phenomenon has been observed across multiple question-answering generation tasks Li et al.
-(2024a); Zheng et al.
-(2022), and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Ignoring input errors (Figure 6 Q.1).
+This phenomenon has been observed across multiple question-answering generation tasks [33, 30], and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Ignoring input errors (Figure 6 Q.1).
 The model ignores the input errors and gives an answer, resulting in the model assigning the same weight to each context given by the user, which in turn affects the generation of the answer.
@@
 To address inconsistent instruction reasoning issues, existing solutions primarily adopt the following two approaches:
-Knowledge Updating.
-SituatedQA Zhang and Choi (2021) attempts to enhance model performance by updating the knowledge base.
-Additionally, CDConv Zheng et al.
-(2022) simulates common user behaviors to trigger chatbots through an automated dialogue generation method, generating contradictions for training purposes.
-Confidence Calibration.
+Knowledge Updating
+SituatedQA [29] attempts to enhance model performance by updating the knowledge base.
+ContraDoc [33] trains models to identify contradictions in long texts using human-annotated datasets.
+Additionally, CDConv [30] simulates common user behaviors to trigger chatbots through an automated dialogue generation method, generating contradictions for training purposes.See Figure 3 for more methods.
+Confidence Calibration
 Given the high cost of data annotation and model fine-tuning, some researchers have sought alternative approaches by introducing additional processing techniques.
-CD2 Jin et al.
-(2024) maximizes probabilistic output and calibrates model confidence under knowledge conflicts using conflict decoupling and comparison decoding methods.
+CD2 [34] maximizes probabilistic output and calibrates model confidence under knowledge conflicts using conflict decoupling and comparison decoding methods.
+MacNoise [39] enhances contradiction retrieval in an augmented generative system by explicitly fine-tuning a discriminator or prompting LLMs to improve contradiction discrimination.
+See Figure 3 for more methods.
+Figure 7: Case of Misinformation Reasoning3.2), caused by temporal misalignment leading to failure responses (Q.1) or data contamination resulting in misleading outputs (Q.2).
 3.2 Misinformation Reasoning
-Erroneous instructions mislead model outputs more severely than inconsistent ones, as they lack obvious contradictions, requiring the model to comprehend, reason, compare input knowledge with its parameterized knowledge, and make objective judgments Cheang et al.
-(2023); Xu et al.
-(2024).
+Erroneous instructions mislead model outputs more severely than inconsistent ones, as they lack obvious contradictions, requiring the model to comprehend, reason, compare input knowledge with its parameterized knowledge, and make objective judgments [102, 103].
 From the input perspective, this paper classifies the sources of erroneous information into two categories as follows: 1) Temporal Alignment Failure (Figure 7 Q.1).
@@
 To solve the above problems, existing methods mainly focus on improving model susceptibility in the face of internal and external knowledge conflicts through targeted fine-tuning and processing.
-CKL Jang et al.
-(2022) ensures that the model’s knowledge is updated in a timely manner through an online approach, although this approach is slightly weaker than re-training in terms of effectiveness Liska et al.
-(2022).
-RKC-LLMs Wang et al.
-(2023) allows a large model to recognize knowledge conflicts by means of instructional fine-tuning that identifying specific passages of conflicting information.
-BIPIA Yi et al.
-(2025) used adversarial training to combat the effects of information pollution and improve model robustness.
-CAR Weller et al.
-(2024) achieved nearly 20% improvement by discriminating external knowledge that may not be contaminated in the RAG system.
+CKL [40] ensures that the model’s knowledge is updated in a timely manner through an online approach, although this approach is slightly weaker than re-training in terms of effectiveness [41].
+RKC-LLMs [46] allows a large model to recognize knowledge conflicts by means of instructional fine-tuning that identifying specific passages of conflicting information.
+BIPIA [42] used adversarial training to combat the effects of information pollution and improve model robustness.
+CAR [47] achieved nearly 20% improvement by discriminating external knowledge that may not be contaminated in the RAG system.
+See Figure 3 for more methods.
+Figure 8: Case of Fuzzy Language Interpretation3.3), where the model relies on biases for fuzzy queries (C.1) or defaults to a response without seeking clarification (C.2)
 3.3 Fuzzy Language Interpretation
 When user instructions contain fuzzy terms (e.g., polysemy or vagueness), LLMs may select an incorrect interpretation from multiple possibilities, potentially leading to misleading responses.
-This phenomenon has been observed across multiple information-seeking tasks Kim et al.
-(2024), and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Self-defined problem (Figure 8 Q.1).
+This phenomenon has been observed across multiple information-seeking tasks [50], and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Self-defined problem (Figure 8 Q.1).
 When the user inputs content with fuzzy sentences, LLMs may choose to generate content based on the preferences of its own training data.
@@
 To solve this problem, researchers have started to parse the semantic information expressed by users through clue engineering.
-Folkscope Yu et al.
-(2023) proposed the FolkScope framework, which uses a large language model to analyze and discriminate users’ fuzzy purchasing intention.
-Miko Lu et al.
-(2024) introduces a hierarchical intention generation framework that interprets users’ posting behaviors by analyzing the fuzzy information they share on social platforms.
-ATC Deng et al.
-(2023) utilizes the Active Thinking Chain cueing scheme, which enhances the proactivity of a biglanguage model by adding goal-planning capabilities to the descriptive reasoning chain.
+Folkscope [48] proposed the FolkScope framework, which uses a large language model to analyze and discriminate users’ fuzzy purchasing intention.
+Miko [49] introduces a hierarchical intention generation framework that interprets users’ posting behaviors by analyzing the fuzzy information they share on social platforms.
+Zhang et al.
+[53] have employed behavioral cloning by using demonstration data from a strong model to train a weaker model so that the weaker model can perform better on a similar task.
+ATC [54] utilizes the Active Thinking Chain cueing scheme, which enhances the proactivity of a biglanguage model by adding goal-planning capabilities to the descriptive reasoning chain.
+See Figure 3 for more methods.
+Figure 9: Case of Intention Clarification Failure3.4) , where it misinterprets sarcasm (Q.1) or ignores prior emotional context (Q.2)
 3.4 Intention Clarification Failure
@@
 2) Ignores prior emotional context (Figure 9 Q.2), when LLMs focus only on the second half of the sentence and ignore the emotions in the previous round of dialog.
-To solve above problems, researchers have started to try to construct multi-domain datasets Chen et al.
-(2023) containing implicit intentions to strengthen the ability of the LLMs to reason about complex intentions and user emotions in multi-round interaction scenarios.
-DeepSeek-R1 Guo et al.
-(2025) enhances its understanding of human intention through a structured process with two RL stages for refining reasoning patterns and aligning with human preferences.
-S1 Muennighoff et al.
-(2025) uses budget forcing to control the number of thinking tokens.
-The upper limit is terminated early by a delimiter, while the lower limit prohibits delimiters and adds "wait" to guide in-depth reasoning and optimize the quality of the answer.
-MoChat Mo et al.
-(2024) constructs multi-round dialogues for spatial localization by using joint grouping spatio-temporal localization.
-Cat. Benchmark Year Lang. Num. Type Description
-Instrution Understanding BotChat Duan et al. (2024) 2023 En&Zh 7658 M Multi-round dialogue eval. via simulated data
-MINT Wang et al. (2024) 2023 En 29,307 M Tool use and feedback in multi-turn dialogue
-MT-BENCH-101 Bai et al. (2024) 2024 En 1,388 M Multi-turn dialogue ability
-$\infty${B}ench Zhang et al. (2024c) 2024 En&Zh 100k S Long-context handling
-L-Eval An et al. (2024) 2024 En 2,000 S Evaluation of Long-Context Language Models
-LongICLBench Li et al. (2025) 2025 En 2,100k S Long In-Context Learning
-Intention Reasoning BIPIA Yi et al. (2025) 2023 En 712.5K S Vulnerability to hint injection
-Miko Lu et al. (2024) 2024 En 10k S Multimodal social intent understanding
-CONTRADOC Li et al. (2024a) 2023 En 891 S Self-contradiction in long docs
-CDCONV Zheng et al. (2022) 2022 Zh 12K M Contradiction in Chinese dialogues
-Relibale Dialog Generate Open-LLM-Leaderboard Ye et al. (2024) 2024 En 10K S Uncertainty in generation
-ETHICS Hendrycks et al. (2021) 2020 En 130K S Moral reasoning
-FACTOR Muhlgay et al. (2024) 2023 En 300 S Factuality in generated text
-Table 1: A selection of widely used benchmark datasets for evaluating LLMs. Th“Cat.”: task stage; ’Lang.’: language of the benchmark, a; ’Num.’: data size; ’Type’: S=Single-round, M=Multi-round.
+To solve above problems, researchers have started to try to construct multi-domain datasets [57] containing implicit intentions to strengthen the ability of the LLMs to reason about complex intentions and user emotions in multi-round interaction scenarios.
+DeepSeek-R1 [55] enhances its understanding of human intention through a structured process with two RL stages for refining reasoning patterns and aligning with human preferences.
+S1 [56] uses budget forcing to control the number of thinking tokens.
+The upper limit is terminated early by a delimiter, while the lower limit prohibits delimiters and adds ”wait” to guide in-depth reasoning and optimize the quality of the answer.
+SoulChat [57] fine-tuned LLMs by constructing a dataset containing more than 2 million samples of multi-round empathic conversations.
+MoChat [58] constructs multi-round dialogues for spatial localization by using joint grouping spatio-temporal localization.
+LARA [68] combines a fine-tuned smaller model with retrieval enhancement mechanisms and integrates it into the architecture of LLM.
+See Figure 3 for more methods.
 4 Reliable Dialog Generation
@@
 While effective on familiar data, they often produce unstable or incomplete responses to unseen inputs, undermining reliability.
+Figure 10: Case of Unstable Content Generation (§4.1), where the model fabricates details when it lacks relevant knowledge (Q.1) or produces incorrect contextual information despite having relevant knowledge (Q.2)
 4.1 Response Stability
 The knowledge acquired by the LLMs is generally determined in the pre-training stage and stored in a parameterized form.
-For data in a specific field, the current model is generally optimized by fine-tuning instructions so that it outputs what humans want Radford et al.
-(2019).
+For data in a specific field, the current model is generally optimized by fine-tuning instructions so that it outputs what humans want [104].
 If knowledge samples that the model has not seen are used in instruction-tuning, it will inevitably cause the model to give a definite response to unknown inputs, and there is a high probability that an answer will be fabricated.
@@
 To address these issues, researchers have explored uncertainty, which quantifies the credibility and stability of model outputs.
-Fine-tuning LLMs.
-To make LLMs more accurate in estimating uncertainty, existing methods fine-tune models Sensoy et al.
-(2018); Amini et al.
-(2020).
-LUQ Zhang et al.
-(2024a) is a novel sampling-based uncertainty quantification method specifically designed for long texts.
-ConformalFactuality Mohri and Hashimoto (2024) defines the associated uncertainties for each possible output.
-External Tools.
+Fine-tuning LLMs
+To make LLMs more accurate in estimating uncertainty, existing methods fine-tune models [69, 70].
+LUQ [80] is a novel sampling-based uncertainty quantification method specifically designed for long texts.
+UaIT [79] using semantic entropy to assess output uncertainty.
+ConformalFactuality [71] defines the associated uncertainties for each possible output.
+See Figure 3 for more methods.
+External Tools
 Fine-tuning LLMs typically demands substantial computing resources and slow training; therefore, reducing computational overhead is crucial for improving efficiency.
-Researcher has proposed methods to evaluate the uncertainty of model outputs through external tools Liu et al.
-(2024).
-ConfidenceElicitation Xiong et al.
-(2024) is a new uncertainty measurement tool for large model outputs.
-CalibrateMath Lin et al.
-assesses uncertainty by requiring models to generate numerical answers with confidence levels, evaluating their reliability.
+Researcher has proposed methods to evaluate the uncertainty of model outputs through external tools [81].
+ConfidenceElicitation [105] is a new uncertainty measurement tool for large model outputs.
+CalibrateMath [82] assesses uncertainty by requiring models to generate numerical answers with confidence levels, evaluating their reliability.
+Figure 11: Case of Misalignment with Human Values (§4.2), where the model generates harmful or offensive content (Q.1)
 4.2 Alignment
 Despite the impressive capabilities of large language models (LLMs), they have raised significant concerns regarding the unsafe or harmful content they may generate.
-LLMs are typically trained on vast datasets scraped from the internet, including inappropriate or harmful content Bender et al.
-(2021).
+LLMs are typically trained on vast datasets scraped from the internet, including inappropriate or harmful content [84].
 This means that the models may inadvertently produce outputs misaligned with human values as follows: 1) Generation of Toxic Content (Figure 11 Q.1).
-LLMs may generate toxic content, such as hate speech or offensive comments, when asked to respond to sensitive topics Luong et al.
-(2024); Dutta et al.
-(2024).
+LLMs may generate toxic content, such as hate speech or offensive comments, when asked to respond to sensitive topics [106, 107].
 2) Conflicts with Moral/Ethical Standards (Figure 11 Q.1).
-LLMs might produce outputs that conflict with moral or ethical standards, such as guiding illegal activities Ramezani and Xu (2023); Abdulhai et al.
-(2024).
+LLMs might produce outputs that conflict with moral or ethical standards, such as guiding illegal activities [108, 109].
 To tackle the above concerns regarding unsafe or harmful content produced by LLMs, researchers have focused on various stages:
-Pretraining Data Cleaning and Curation.
-To minimize the risks associated with harmful or inappropriate content, LLM training datasets should undergo rigorous cleaning processes Bender et al.
-(2021), such as filtering out toxic language, hate speech, and harmful stereotypes.
-Tools like word embedding debiasing methods Rakshit et al.
-can help identify and remove toxic and biased content.
-Reinforcement Learning-based Alignment.
-To further align LLMs with human and societal norms, reinforcement learning methods, such as RLHF and its advanced variants, such as PPO Ouyang et al.
-(2022), DPO Zeng et al.
-(2024), and GRPO Shao et al.
-(2024), are very essential.
-As an extension of RLHF, RLAIF Lee et al.
-(2024) leverages AI systems to assist in the feedback process, making evaluation and fine-tuning more scalable.
-In-context Alignment.
+Pretraining Data Cleaning and Curation
+To minimize the risks associated with harmful or inappropriate content, LLM training datasets should undergo rigorous cleaning processes [84], such as filtering out toxic language, hate speech, and harmful stereotypes.
+Tools like the Perspective API [83] and word embedding debiasing methods [85] can help identify and remove toxic and biased content.
+Tools like word embedding debiasing methods [85] can help identify and remove toxic and biased content.
+Reinforcement Learning-based Alignment
+To better align LLMs with human values and societal norms, reinforcement learning approaches are widely adopted, including RLHF and its advanced variants such as PPO [3], DPO [86], and GRPO [87].
+Extending RLHF, RLAIF [88] leverages AI systems to assist in the feedback process, improving the scalability of evaluation and fine-tuning while enabling continuous monitoring and refinement of model behavior [88].
+In-context Alignment
 It leverages the ability of LLMs to adapt their responses based on a few examples provided in the prompt.
-Lin et al.
-(2024) demonstrates that effective alignment can be achieved purely through ICL with just a few stylistic examples and a system prompt.
-Huang et al.
-(2024a) explored the effectiveness of different components of In-context alignment, and found that examples within the context are crucial for enhancing alignment capabilities.
-5 Case Analysis
-To demonstrate the challenges of existing LLMs in above stages, we provide cases analysis in Appendix B, and make statistical comparisons in Table 2.
-The results show that models with reasoning capabilities (such as GPT-o3*, Deepseek-R1*) perform better in command understanding tasks such as long text understanding and multi-round dialogue, but have not completely solved all problems.
-In the intention reasoning task, except for "Inconsistent information reasoning", the models generally performed poorly, and only a few models partially passed the case.
-In terms of reliable dialogue generation, all models performed poorly in response stability, and only GPT-o3 and Deepseek-R1 performed well in the alignment test.
-Overall, while reasoning abilities contribute to improved task performance, significant limitations remain across all three stages.
-6 Benchmark
-This section covers benchmarks for LLMs in above three stage (Table 1), more details see Appendix A.
-6.1 Benchmarking Instruction Understanding
-Instruction understanding in LLMs involves extracting key information, maintaining coherence, and adapting to dynamic conversation changes, especially in long or multi-round dialogues.
-LLM capabilities in multi-turn dialogue and long-context processing have been explored through various benchmarks.
-BotChat Duan et al.
-(2024) evaluates dialogue generation, showing GPT-4’s strengths but noting instruction compliance and length limitations in other models.
-MINT Wang et al.
-(2024) highlights limited progress in tool use and feedback for complex tasks.
-MT-Bench-101 Bai et al.
-(2024) identifies challenges in enhancing long-term interaction skills.
-For long-context tasks, performance drops in ultra-long texts (Zhang et al.
-Zhang et al.
-(2024c)), L-Eval An et al.
-(2024) emphasizes length-instruction-enhanced metrics, and LongICLBench Li et al.
-(2025) reveals difficulties in reasoning across multiple pieces of information.
+[89] demonstrates that effective alignment can be achieved purely through ICL with just a few stylistic examples and a system prompt.
+[90] explored the effectiveness of different components of In-context alignment, and found that examples within the context are crucial for enhancing alignment capabilities.
+Liu et al.
+[91] introduced PICA, which uses a two-stage approach to improve alignment efficiency.
 Stage Challenge GPT-4o GPT-o3* Qwen3 Qwen3* Deepseek-v3 Deepseek-R1*
@@
 ’A’ indicates the model’s accuracy is greater than 75%, ’B’ is between 50% and 75%, and ’C’ is below 50%.
-6.2 Benchmarking LLM Reasoning
-LLM intention reasoning involves inferring user intentions by interpreting both explicit and implicit language cues.
-The existing benchmarks comprehensively evaluate the multifaceted reasoning capabilities of LLMs, encompassing vulnerability to indirect hint injection attacks (BIPIA Yi et al.
-(2025)), advantages in multimodal intention understanding (Miko Lu et al.
-(2024)), analysis of self-contradictions in long documents (CONTRADOC Li et al.
-(2024a)), and contradiction detection in Chinese dialogues (CDCONV Zheng et al.
-(2022)).
-Collectively, these benchmarks highlight both the challenges and advancements of LLMs in complex reasoning.
-6.3 Benchmarking LLM Generation
-LLM Generation assesses a model’s ability to understand user instructions, avoid fabricating false information, and generate accurate, contextually appropriate responses.
-Open-LLM-Leaderboard Ye et al.
-(2024), ETHICS Hendrycks et al.
-(2021) and FACTOR Muhlgay et al.
-(2024) all focus on the reliability evaluation of content generated by large models.
-Open-LLM-Leaderboard finds that large-scale models have higher uncertainty, and fine-tuned models have higher accuracy but greater uncertainty.
-ETHICS focuses on the ethical value alignment of generated content.
-FACTOR evaluates factuality through scalable methods to ensure that diverse and rare facts are covered.
-7 Future Directions
-This section summarizes ongoing challenges in above stages with LLMs and outlines potential future research directions.
-Automated Annotation Framework.
-Although LLMs excel in general-domain tasks, they often produce hallucinated or incomplete content in specialized fields due to limited domain-specific training data.
-While contextual learning and instruction fine-tuning methods have been explored to address this issue, manual data annotation remains labor-intensive and prone to quality inconsistencies.
-An automated annotation framework could streamline data labeling, enhancing model performance in specialized fields by ensuring higher quality and scalability of domain-specific training datasets.
-GraphRAG.
-LLMs have shown impressive language generation capabilities through pre-training on large datasets, but their reliance on static data often results in inaccurate or fictional content, particularly in domain-specific tasks.
-The Graph-enhanced generation approach aims to tackle this by leveraging KGs and GNNs for precise knowledge retrieval.
-Despite its advantages, GraphRAG faces challenges in capturing structural information during graph reasoning tasks and struggles with multi-hop retrieval accuracy and conflict resolution between external and internal knowledge.
-Future work should focus on refining retrieval strategies and improving the stability and accuracy of GraphRAG in complex tasks.
-Balancing Safety and Performance.
-Although advancements in alignment techniques have improved factual accuracy and safety, they often come at the cost of the model’s creativity and fluency.
-Striking a balance between safety and performance is crucial.
-Future research should explore new alignment methods that ensure both the safety and usability of LLMs, optimizing the trade-off between generating reliable, safe content and maintaining the model’s creative and contextual capabilities.
-8 Conclusion
-This paper analyzes LLMs’ performance in processing user instructions.
-Despite progress in natural language understanding, LLMs struggle with complex, inconsistent instructions, often resulting in biases, errors, and hallucinations.
-Improvements through prompt engineering, model expansion, and RLHF et al.
-have not fully addressed LLMs’ limitations in reasoning and comprehension, limiting their real-world applicability.
-We identify three challenges: instruction understanding, intention reasoning and reliable dialog generation.
-Future research should focus on enhancing reasoning for complex instructions and aligning outputs with user intention to improve LLMs’ reliability.
-Limitations
-We have made significant efforts to ensure the quality of this study, but certain limitations may still exist.
-First, due to constraints on the length of the discussion, this article primarily focuses on analyzing examples of human interaction with large models, which inevitably limits the depth of methodological details we can explore.
-Second, while our meta-analysis draws extensively from prominent academic platforms such as the ACL conference series, ICLR, ICML, and the arXiv preprint repository, it is possible that some valuable insights from other sources have not been included.
-It is worth noting that several open scientific questions in this field remain unresolved, and the academic community has yet to reach a consensus on these issues.
-To address these limitations, we plan to establish a long-term tracking mechanism to monitor new developments in the field.
-This will allow us to incorporate emerging theoretical advancements and dynamically refine or expand upon the perspectives presented in this study.
-Appendix A Benchmark Details
-A.1 Instruction Understanding Benchmark
-BotChat Duan et al.
-(2024) specifically addresses the evaluation of LLMs’ ability to emulate human-like, multi-turn conversations using an LLM-centric approach.
-This benchmark assesses utterance generation, evaluation protocols, and judgment.
-Findings suggest that models like GPT-4 demonstrate exceptional performance as both generators and judges, producing human-indistinguishable dialogues and showing high alignment with human evaluation standards.
-Conversely, other LLMs face challenges in generating quality multi-turn dialogues due to issues like inadequate instruction-following and a tendency towards prolixity, particularly in generating extensive dialogues.
-MINT Wang et al.
-(2024) introduces a benchmark that evaluates LLMs’ capacity to solve complex tasks through multi-turn interactions, emphasizing the use of external tools and leveraging natural language feedback.
-MINT provides a reproducible evaluation framework where LLMs can access tools and receive simulated user feedback.
-Analysis using MINT reveals that LLMs generally benefit from tools and language feedback.
-Interestingly, studies with MINT suggest that techniques like supervised instruction-finetuning (SIFT) and reinforcement learning from human feedback (RLHF) might not consistently improve multi-turn capabilities, indicating the need for further research in training methodologies for conversational settings.
-MT-BENCH-101 Bai et al.
-(2024) offers a structured approach with its three-layer hierarchical capability taxonomy and a dataset of 1,388 dialogue pairs across 13 tasks.
-This allows for a fine-grained evaluation of LLMs’ multi-turn dialogue skills.
-The findings from studies utilizing MT-BENCH-101 indicate that commonly employed techniques yield limited improvements in multi-turn performance, suggesting a need for more effective methods to enhance conversational abilities over extended interactions.
-Zhang et al.
-Zhang et al.
-(2024c) focus on the challenge of long-context processing, introducing a benchmark of 12 tasks, each averaging over 100K tokens, to assess LLMs’ long-context processing abilities.
-Results indicate a marked performance drop with longer contexts, highlighting the need for further improvement.
-L-Eval An et al.
-(2024) contributes to standardizing the evaluation of Long-Context Language Models (LCLMs) by providing a new evaluation suite with diverse tasks, document lengths (3k to 200k tokens), and a large number of human-labeled query-response pairs.
-It also investigates the effectiveness of evaluation metrics, suggesting that Length-instruction-enhanced (LIE) evaluation and LLM judges correlate better with human judgments.
-Comprehensive studies using L-Eval offer insights into the performance of both commercial and open-source LCLMs.
-LongICLBench Li et al.
-(2025) focuses on long in-context learning in extreme-label classification, with input lengths ranging from 2K to 50K tokens and a large number of classes.
-Evaluation on this benchmark showed that while LLMs perform well on less challenging tasks, they struggle with more complex ones involving a large label space and longer contexts, revealing a bias towards later parts of the input and difficulty in reasoning over multiple pieces of information.
-A.2 LLM Intention Reasoning Benchmark
-BIPIA Yi et al.
-(2025) evaluates LLMs under indirect hint injection attacks across five scenarios and 250 targets, revealing vulnerabilities in all models, with GPT-3.5-turbo and GPT-4 exhibiting notably higher susceptibilities.
-Miko Lu et al.
-(2024) assesses multimodal models in understanding social media user intentions.
-The benchmark, which includes 979 social media entries, shows that multimodal LLMs outperform text-only models like LLama2-7B and GLM4.
-Incorporating image data enhances the model’s ability to interpret user intentions, improving accuracy.
-CONTRADOC Li et al.
-(2024a) is the first dataset for analyzing self-contradictions in long documents.
-Evaluation of GPT-4, PaLM2, and other LLMs on this dataset reveals that while GPT-4 outperforms others and even surpasses human performance, it still struggles with complex contradictions requiring nuanced reasoning.
-CDCONV Zheng et al.
-(2022) focuses on contradiction detection in Chinese dialogues, containing over 12,000 dialogue rounds.
-The study shows that the Hierarchical method consistently outperforms others in detecting contradictions, highlighting the importance of accurate contextual modeling in dialogue understanding.
-A.3 LLM Reliable Generation Benchmark
-Open-LLM-Leaderboard Ye et al.
-(2024) introduces a novel benchmark that integrates uncertainty quantification to evaluate the reliability of content generation across tasks like QA, comprehension, and dialogue.
-Results show that larger models often exhibit greater uncertainty, and fine-tuned models tend to have higher uncertainty despite higher accuracy.
-ETHICS Hendrycks et al.
-(2021) evaluates whether generated content aligns with human ethical values, such as justice and well-being.
-The study finds that while models like GPT-3 show promise in predicting human moral judgments, they still need improvement in this domain.
-FACTOR Muhlgay et al.
-(2024) addresses the evaluation of factuality in LLMs by providing a scalable method that ensures diverse and rare facts are considered.
-Testing models such as GPT-2 and GPT-Neo show that, while benchmark scores correlate with perplexity, they better reflect factuality in open-ended generation, especially when retrieval augmentation is applied.
-Appendix B Challenges Cases in Human-LLMs Aligement
-To fully understand the limitations of large language models (LLMs) in practical applications, it is particularly important to analyze their performance in a variety of complex scenarios.
-We collected human user instructions from real scenarios that the large model needs to interact with on a total of eight challenges that the summarized existing large language models (LLMs) face in the three phases of instruction understanding, intention reasoning, and reliable dialog generation, and cleaned and filtered the instructions through manual screening, diversity sampling, and difficulty filtering, and finally, 50 human instructions were used for each challenge.
+5 Challenges Cases in Human-LLMs Aligement
+To fully understand the limitations of LLMs in practical applications, it is particularly important to analyze their performance in a variety of complex scenarios.
+We collected human user instructions from real scenarios that the large model needs to interact with on a total of eight challenges that the summarized existing LLMs face in the three phases of instruction understanding, intention reasoning, and reliable dialog generation, and cleaned and filtered the instructions through manual screening, diversity sampling, and difficulty filtering, and finally, 50 human instructions were used for each challenge.
 The statistical results of the test are shown in Table.1.
@@
 Through these cases, we systematically reveal the typical failure modes of each model in different aspects and their deep-seated causes, providing important references and directions for targeted optimization in future research.
-B.1 Case of Instruction Understanding
-Although large language models (LLMs) perform well in short text and single-round dialogue scenarios, their ability to understand and execute complex instructions in long contexts and multi-round dialogues still faces many severe challenges.
+5.1 Case of Instruction Understanding
+Although LLMs perform well in short text and single-round dialogue scenarios, their ability to understand and execute complex instructions in long contexts and multi-round dialogues still faces many severe challenges.
 Figure 4 shows that long texts usually contain a lot of irrelevant or redundant information, which causes the core content directly related to the task to be submerged.
@@
 In multi-round dialogue scenarios, the model is not only prone to gradually accumulate errors due to inaccurate understanding of the previous text, but may also fail to correctly judge the relevance of each round of dialogue, causing the answer to deviate from the user’s real needs.
-As shown in Table 2, Figure 4, and Figure 5, in the instruction understanding stage, the reasoning models generally performed well on the "Long-Text Comprehension" task, while the non-reasoning models all struggled, and for “Multi-Turn Conversation Handling”, in addition to the non-reasoning models, Qwen3, which provides reasoning capability, also exhibited suboptimal performance.
+As shown in Table 1, Figure 4, and Figure 5, in the instruction understanding stage, the reasoning models generally performed well on the ”Long-Text Comprehension” task, while the non-reasoning models all struggled, and for “Multi-Turn Conversation Handling”, in addition to the non-reasoning models, Qwen3, which provides reasoning capability, also exhibited suboptimal performance.
 The above results show that reasoning can effectively improve the ability of LLMs in instruction understanding, but cannot completely solve the instruction understanding problem.
-B.2 Case of Intention Reasoning
+5.2 Case of Intention Reasoning
 Since there are often spelling errors, factual contradictions and semantic ambiguities in user instructions, LLM faces many challenges in understanding the user’s true intentions.
@@
 This suggests that all models have significant challenges in reasoning about intentions.
-B.3 Case of Reliable Dialog Generation
+5.3 Case of Reliable Dialog Generation
 Current large language models exhibit both deterministic response preferences and ambiguous knowledge boundaries during generation.
@@
 Overall, reasoning capabilities can improve the performance of llms for interacting with real human user commands, however, all models still face significant challenges and remain underpowered in real-world and human interaction scenarios.
-Figure 4: Case of Remote Information Failure (§2.1), where the model forgets relevant information over long distances in long context.
-Figure 5: Case of Incorrect Relevance Judgment2.2), such as the model incorrectly associates wrong content from the previous turn.
-Figure 6: Case of Inconsistent Instruction Reasoning (§3.1), where LLMs fail to detect conflicting inputs (Q.1) or overlooking logical inconsistencies (Q.2)
-Figure 7: Case of Misinformation Reasoning3.2), caused by temporal misalignment leading to failure responses (Q.1) or data contamination resulting in misleading outputs (Q.2).
-Figure 8: Case of Fuzzy Language Interpretation3.3), where the model relies on biases for fuzzy queries (C.1) or defaults to a response without seeking clarification (C.2)
-Figure 9: Case of Intention Clarification Failure3.4) , where it misinterprets sarcasm (Q.1) or ignores prior emotional context (Q.2)
-Figure 10: Case of Unstable Content Generation (§4.1), where the model fabricates details when it lacks relevant knowledge (Q.1) or produces incorrect contextual information despite having relevant knowledge (Q.2)
-Figure 11: Case of Misalignment with Human Values (§4.2), where the model generates harmful or offensive content (Q.1)
+Cat. Benchmark Year Lang. Num. Type Description
+Instrution Understanding BotChat [110] 2023 En&Zh 7658 M Multi-round dialogue eval. via simulated data
+MINT [97] 2023 En 29,307 M Tool use and feedback in multi-turn dialogue
+MT-BENCH-101 [111] 2024 En 1,388 M Multi-turn dialogue ability
+$\infty${B}ench [112] 2024 En&Zh 100k S Long-context handling
+L-Eval [113] 2024 En 2,000 S Evaluation of Long-Context Language Models
+LongICLBench [114] 2025 En 2,100k S Long In-Context Learning
+Intention Reasoning BIPIA [42] 2023 En 712.5K S Vulnerability to hint injection
+Miko [49] 2024 En 10k S Multimodal social intent understanding
+CONTRADOC [33] 2023 En 891 S Self-contradiction in long docs
+CDCONV [30] 2022 Zh 12K M Contradiction in Chinese dialogues
+Relibale Dialog Generate Open-LLM-Leaderboard [115] 2024 En 10K S Uncertainty in generation
+ETHICS [116] 2020 En 130K S Moral reasoning
+FACTOR [117] 2023 En 300 S Factuality in generated text
+Table 2: A selection of widely used benchmark datasets for evaluating LLMs. Th“Cat.”: task stage; ’Lang.’: language of the benchmark, a; ’Num.’: data size; ’Type’: S=Single-round, M=Multi-round.
+6 Benchmark
+This section covers benchmarks for LLMs in above three stage (Table 2).
+6.1 Benchmarking Instruction Understanding
+Instruction understanding in LLMs involves extracting key information, maintaining coherence, and adapting to dynamic conversation changes, especially in long or multi-round dialogues.
+LLM capabilities in multi-turn dialogue and long-context processing have been explored through various benchmarks.
+BotChat [110] evaluates dialogue generation, showing GPT-4’s strengths but noting instruction compliance and length limitations in other models.
+MINT [97] highlights limited progress in tool use and feedback for complex tasks.
+MT-Bench-101 [111] identifies challenges in enhancing long-term interaction skills.
+For long-context tasks, performance drops in ultra-long texts (Zhang et al.
+[112]), L-Eval [113] emphasizes length-instruction-enhanced metrics, and LongICLBench [114] reveals difficulties in reasoning across multiple pieces of information.
+6.2 Benchmarking LLM Reasoning
+LLM intention reasoning involves inferring user intentions by interpreting both explicit and implicit language cues.
+The existing benchmarks comprehensively evaluate the multifaceted reasoning capabilities of LLMs, encompassing vulnerability to indirect hint injection attacks (BIPIA [42]), advantages in multimodal intention understanding (Miko [49]), analysis of self-contradictions in long documents (CONTRADOC [33]), and contradiction detection in Chinese dialogues (CDCONV [30]).
+Collectively, these benchmarks highlight both the challenges and advancements of LLMs in complex reasoning.
+6.3 Benchmarking LLM Generation
+LLM Generation assesses a model’s ability to understand user instructions, avoid fabricating false information, and generate accurate, contextually appropriate responses.
+Open-LLM-Leaderboard [115], ETHICS [116] and FACTOR [117] all focus on the reliability evaluation of content generated by large models.
+Open-LLM-Leaderboard finds that large-scale models have higher uncertainty, and fine-tuned models have higher accuracy but greater uncertainty.
+ETHICS focuses on the ethical value alignment of generated content.
+FACTOR evaluates factuality through scalable methods to ensure that diverse and rare facts are covered.
+7 Future Directions
+This section summarizes ongoing challenges in above stages with LLMs and outlines potential future research directions.
+Automated Annotation Framework
+Although LLMs excel in general-domain tasks, they often produce hallucinated or incomplete content in specialized fields due to limited domain-specific training data.
+While contextual learning and instruction fine-tuning methods have been explored to address this issue, manual data annotation remains labor-intensive and prone to quality inconsistencies.
+An automated annotation framework could streamline data labeling, enhancing model performance in specialized fields by ensuring higher quality and scalability of domain-specific training datasets.
+GraphRAG
+LLMs have shown impressive language generation capabilities through pre-training on large datasets, but their reliance on static data often results in inaccurate or fictional content, particularly in domain-specific tasks.
+The Graph-enhanced generation approach aims to tackle this by leveraging KGs and GNNs for precise knowledge retrieval.
+Despite its advantages, GraphRAG faces challenges in capturing structural information during graph reasoning tasks and struggles with multi-hop retrieval accuracy and conflict resolution between external and internal knowledge.
+Future work should focus on refining retrieval strategies and improving the stability and accuracy of GraphRAG in complex tasks.
+Quantifying Uncertainty in LLMs
+As LLMs like LLama and ChatGPT have revolutionized content generation, ensuring the reliability of their outputs remains a critical challenge.
+Uncertainty quantification is a promising approach to address this, enabling models to provide a confidence assessment alongside their responses.
+Evidence learning [69, 70], an emerging method in uncertainty representation, offers a reliable approach for quantifying uncertainty directly from data.
+However, its application to LLMs is computationally intensive, and most existing work focuses on small-scale models.
+Future research should aim to optimize uncertainty quantification for large-scale models efficiently.
+Balancing Safety and Performance
+Although advancements in alignment techniques have improved factual accuracy and safety, they often come at the cost of the model’s creativity and fluency.
+Striking a balance between safety and performance is crucial.
+Future research should explore new alignment methods that ensure both the safety and usability of LLMs, optimizing the trade-off between generating reliable, safe content and maintaining the model’s creative and contextual capabilities.
+8 Conclusion
+This paper analyzes LLMs’ performance in processing user instructions.
+Despite progress in natural language understanding, LLMs struggle with complex, inconsistent instructions, often resulting in biases, errors, and hallucinations.
+Improvements through prompt engineering, model expansion, and RLHF et al.
+have not fully addressed LLMs’ limitations in reasoning and comprehension, limiting their real-world applicability.
+We identify three challenges: instruction understanding, intention reasoning and reliable dialog generation.
+Future research should focus on enhancing reasoning for complex instructions and aligning outputs with user intention to improve LLMs’ reliability.